### v1

# Airdrops: Giving Money Away Is Harder Than It Seems

Abstract.
Airdrops are used by blockchain applications and platforms to attract an initial user base, and to grow the user base over time. In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying platform, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop has been completed. Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
In this work, we outline the design space for airdrops, and specify a reasonable list of outcomes that an airdrop should ideally result in. We then analyze on-chain data from several larger-scale airdrops to empirically evaluate the success of previous airdrops, with respect to our desiderata. In our analysis, we demonstrate that airdrop farmers frequently dispose of the lion’s share of airdrops proceeds via exchanges. Our analysis is followed by an overview of common pitfalls that common airdrop designs lend themselves to, which are then used to suggest concrete guidelines for better airdrops.
1. Introduction
Blockchain platforms commonly create reward schemes to lure new customers and maintain the loyalty of existing users. In recent years, giving platform-minted tokens to users, or airdropping tokens, has proven itself as a particularly popular scheme (Makridis et al., 2023). Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives. Intuitively, this is not optimal, and may result in wasting funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
While the basic premise of airdrops is relatively straightforward, the design space for such reward schemes is large, and may depend on the characteristics of the platform designing the airdrop. For example, some airdrop mechanisms target power-users and endow them with a generous amount of rewards in the hopes that these users will generate valuable economic activity, and thus attract other users as well.
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process. Specifically, we analyze the data pertaining to four recent airdrops and conclude that the majority of proceeds (as high as 95%, typically only within a couple of transfer steps) have been sold via exchanges, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers as opposed to regular users. This comparison is followed by a description of common pitfalls faced by airdrops in the past. Given that the blockchain ecosystem is still quite young, it is to be expected that the theory and practice of blockchain airdrops is still in its infancy, and thus that previous airdrops may have not been entirely successful. Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms which are more fair to honest users.
1.1. Contributions
Arbitrum study. We perform a comprehensive case study of the Arbitrum (Kalodner et al., 2018) airdrop by measuring elements such as transaction volume, token, distribution structure, and token value before and after the airdrop period.
Quantitative analysis. We quantitatively analyze airdrops executed by ENS (ENS, 2023), dYdX (Juliano, 2018), Gemstone (Haig, 2023; lingland09, 2023), and 1inch (Bukov and Melnik, 2020). We show that most of the funds received through these airdrops are sold on exchanges, instead of being used to create dapp or platform engagement on the part of end-users.
Qualitative analysis. We qualitatively analyze previous airdrops, and propose guidelines for developing future airdrops in an effort to avoid some of the pitfalls in the airdrops we analyze.
Scientific reproducibility. To ensure the scientific reproducibility of our empirical analysis, we share our dataset and scripts (Messias et al., 2023b). Refer Section A in the appendix for more details.
2. Airdrop Desiderata
Airdrops are typically used to bootstrap a community of users. We break down this overarching high-level goal into multiple sub-goals. These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
1. Attract users in the short-term. Historically, airdrops were used by new blockchain platforms to create an initial user base. In particular, decentralized platforms commonly become more useful and attractive to users as a function of the amount of economic activity taking place.
2. Foster long-term involvement. Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time. Ideally, platform users should become regular, everyday users. This can be achieved by issuing rewards that can only be used within the platform, similarly to frequent-flyer points. For example, in layer 2 blockchains (L2s) (Sguanci et al., 2021), one may award discounts on future transactions. Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the platform has to offer. For example, Linea’s airdrop quests gave users an in-depth tour of the platform’s various features and use-cases (Labs, 2023a).
3. Target value-creating users. An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of the platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools (Yaish et al., 2022) and decentralized exchanges (DEXs) (Yaish et al., 2023), or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable. Such users provide additional use-cases for a platform, and thus serve to attract other users.
Table 1. Airdrops we analyze in this paper.
Airdrop Start date End date Blockchain
ENS October 2021 May 2022 Ethereum
dYdX September 2021 Still active Ethereum
1inch December 2020 Still active Ethereum
[…]
3. Exchange Aftermarket Analysis
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. Thus, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them (jhackworth, 2022), and 61% have sold their tokens shortly after the drop in the case of ParaSwap (Fan et al., 2023). In both cases, recipients mostly stopped using the corresponding platforms within months. This behavior indicates that the airdrop did not succeed in retaining recipients, or that recipients are mostly sybil accounts. Furthermore, a quick sell-off may alarm the market, if it is interpreted as a sign of a lack of belief in the long-term success of the platform.
Below we consider data related to the four airdrops (ENS, dYdX, Gemstone, and 1inch) in Table 1 that were collected from an Ethereum and zkSync Era archive nodes. We have made available both our data collection and analysis scripts, along with the dataset itself (Messias et al., 2023b). To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query:
select * from addresses_ethereum.dex
Platform Token Tokens % sold at # of recipient Tokens received per address
supply distributed exchanges addresses Mean Std. Median Min Max
ENS $1\times 10^{8}$ $1.96\times 10^{7}$ 36.51% $102822$ $190.93$ $141.57$ $152.4$ $0.0001$ $1143.54$
dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.35% $42682$ $2325.85$ $91336.75$ $392.56$ 00 $2.47\times 10^{7}$
1inch $1.5\times 10^{9}$ $8.3\times 10^{7}$ 53.47% $43402$ $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
Gemstone $5.06\times 10^{16}$ $5.04\times 10^{16}$ 95.01% $20014$ $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
Table 2. Airdrop distribution statistics for the four airdrops.
3.1. Tokens Distributed Over Time
Table 2 proves some numbers on how airdropped tokens are spread across the user base. We observe that the number of recipients is quite high for every analyzed airdrops, which may indicates some effort on the part of airdrop farmers (column 3). Moreover, the data indicates that the tokens were often traded on exchanges, as evident from the fact that the first transfer made by the addresses after the airdrop was a sale in an exchange (column 10). Gemstone stands out from the rest, with 95% of its tokens ending up on exchanges.
Furthermore, Gemstone’s airdrop stands out significantly in terms of the number of tokens distributed compared to other airdrops. This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops. Remarkably, Gemstone allocated the lion’s share, specifically 99.53%, of their total token supply during their airdrop campaign. It is important to highlight that Gemstone was executed primarily as a sybil attack rather than a legitimate airdrop.
Platform Airdrop recipients # of hops
Total # sent to exchanges % sent to exchanges Mean Std. Median Min. Max.
ENS $102822$ $86157$ 83.79% 1.75 1.06 2 1 50
dYdX $42682$ $41320$ 96.81% 2.25 0.59 2 1 18
1inch $43402$ $41913$ 96.57% 1.85 0.64 2 1 7
Gemstone $20014$ $20000$ 99.93% 1 1 1 1 1
Table 3. Airdrop recipients’ interaction with exchanges. Gemstone stands out as an exception with a median of 1 hop, while other protocols show a median of 2 hops for connecting airdrop recipients with exchanges.
(a) Over the number of post-airdrop blocks.
(b) Over the number of post-airdrop days.
Figure 1. Time for airdrop tokens to be transferred to exchanges.
Focusing on Table 3, we computed the shortest path taken from each recipient address of an airdrop to any of the exchange address in our set. Our findings reveal that the transfer of tokens to exchanges typically takes only a few steps, suggesting that airdrop farmers, in general, are not making significant efforts to hide their activities. The outlier is Gemstone where in all cases tokens were sold within only one hop. Surprisingly, the majority of accounts managed to reach an exchange with a relatively low number of intermediating hops. This observation underscores the prominent role that exchanges play within the crypto ecosystem.
(a) ENS.
(b) dYdX.
(c) Gemstone.
(d) 1inch.
Figure 2. Network structure for selling airdropped tokens in 4 airdrops.
Figure 1 shows the CDFs that demonstrate how quickly tokens are sold at changes. Per Figure 0(a), we see that within about a million blocks, the majority of tokens are gone. For Gemstone the number of blocks significantly higher than others because of a delay introduced by airdrop farmers. As the block generation is quite different for Ethereum (a new block is created every 15s on average) and zkSync, we normalized the blocks in terms of the number of days. Per Figure 0(b), within a single day, 64.67% of the 1inch tokens were sent to an exchange. In comparison, ENS had a transfer rate of 45.86%, while dYdX stood at 37.41%. This does not align with one of the primary objectives of a protocol when conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2.
3.2. Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdropped the first few hours after the airdrop and plot their largest component in Figure 2. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer. We show that for all protocols but Gemstone, Uniswap is the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from the addresses receiving the airdrops, followed by SushiSwap.
[…]
(a) Absolute number of unique daily active addresses.
(b) Daily active addresses, relative to the average for the 50 days before Arbitrum’s airdrop.
Figure 3. Unique daily active addresses per platform. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
Figure 4. Median daily transaction fees, in USD.
4. Measuring the Airdrop Lift
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic. In this section, we examine this through the lens of some relevant metrics applied to the Arbitrum airdrop. We rely on data source from Growthepie (Growthepie, 2023) and list details in Section A in appendix.
4.1. Unique Active Addresses
Figure 3 examines the number of unique addresses that were active on each platform using two different approaches. The Figure 2(a) depicts the absolute number of addresses. The Figure 2(b) shows the relative amount of unique daily active addresses for each platform, when compared to the same platform’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 (Rustgi, [n. d.]). This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users (Barthere, 2023).
Observation 1: others outperformed Arbitrum, without an airdrop. The absolute number of addresses has increased for Arbitrum post-airdrop, and remained at least 50% higher than its pre-airdrop level in the four months since then. But, by looking at other platforms, it becomes apparent that a similar result can be attained without an airdrop. This is demonstrated by Optimism’s trajectory in Figure 2(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 (Nijkerk, 2023). This is further exemplified by zkSync Era overtaking Arbitrum’s in Figure 2(a), two months after the airdrop.
Observation 2: fees may explain the closing bap between Arbitrum and Optimism. Arbitrum leads over Optimism with respect to the number of unique daily active addresses. While in the 50 days before the airdrop Arbitrum has enjoyed an average of 2.6 times more unique daily active addresses than Optimism, this gap has recently started to close: the average over the last 50 days in our dataset is considerably lower, and is equal to 1.83 times more. Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 4.
Observation 3: the number of unique addresses oscillates over time. Some of the previous observations exhibit oscillatory behavior, in particular in Figure 2(b) depicting the relative amount of unique addresses: rising, reaching a peak, and then subsequently dropping, all in rapid succession. In particular, the phases of Optimism and Arbitrum are inverted regarding the relative amount of unique addresses. One may attempt to explain this by reasoning that when a peak is reached in one of the platforms, fees should increase due to the load, thus prompting cost-conscious users to turn to a different platform. But, when considering the median transaction fees paid per day (per Figure 4), it is apparent that fees do not have corresponding peaks and dips. In fact, Arbitrum consistently had lower median fees mid-May.
Caveat: the unique address metric is gameable. The overall picture painted so far is that with regard to the daily unique account metric, Arbitrum’s airdrop did not succeed in generating long-lasting user involvement. An important caveat of this metric is that in blockchains, users can typically costlessly create and use multiple addresses. This practice is also called a sybil attack, with attackers also known as sybils. Sybil attacks may benefit users, as some airdrops limit the amount of tokens that a single address can receive. Thus, using the number of unique addresses as a metric may paint an incomplete picture: the large changes we observed may be caused by sybil attackers. Although some employ sophisticated tools to identify sybil addresses, these are heuristic in nature and may fail to capture all of them. Thus, we need to consider other metrics that are hopefully more sybil-resistant and representative of genuine user activity.
4.2. Transaction-related Metrics
A natural next step is to examine transaction-related metrics. In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee (Gafni and Yaish, 2022), when ignoring retroactive rebates or the possibility of platforms operating at a loss (Optimism, 2023).
Observation 4: the Arbitrum-Optimism gap is smaller, when considering transactions. Figure 4(a) depicts the absolute daily transaction count per platform, and Figure 4(b) shows the relative amount of transaction count for each platform compared to the same platform’s average transaction count for the period between February 1st and March 22nd, 2023. These figures are quite similar to the corresponding ones for the unique address count metric.
Some notable differences are that with respect to transaction counts, the gap between Arbitrum and Optimism has nearly closed entirely by the end of July. Furthermore, Figure 4(b) shows that the daily transaction count for Immutable X has almost halved since Arbitrum’s airdrop. Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the platform.
(a) Absolute daily transactions.
(b) Daily transactions relative to the average for the 50 days before Arbitrum’s airdrop.
Figure 5. Daily number of transactions per platform. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
Observation 5: Arbitrum’s per address transaction count dropped since the airdrop. To better understand how Arbitrum’s airdrop affected user engagement in other platforms, we now present the Figure 6, which depicts the relative average daily transaction count per unique address. The figure shows that since the drop, Arbitrum’s per-user transaction count has fallen to less than 75% of its pre-airdrop level.
Figure 6. Daily transactions per address, relative to the average before Arbitrum’s airdrop.
Caveat: transaction counts are inaccurate when ignoring fees. Relying on the previous metrics may be inaccurate: a high transaction count does not necessarily signal high user engagement. In particular, some platforms require users to perform a certain number of transactions over some time frame to become eligible for their airdrops. Thus, if fees are low, a high number of transactions per account may be caused by users who perform economically-useless transactions, for the sole purpose of receiving the drop. This brings to mind Goodhart’s law (Manheim and Garrabrant, 2019), which posits that “when a measure becomes a target, it ceases to be a good measure”.
Observation 6: since June, average transaction fees are similar for all platforms. Conceptually, good metrics should measure a user’s amount of skin in the game. A possible proxy for that is the amount of funds the user is willing to expend for interacting with the platform. When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all platforms are quite similar in both regards. Furthermore, when comparing the average fee per transaction to the median one (per Figures 4 and 7), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
Figure 7. Daily average transaction fee spent per transaction per day, in USD.
Another interesting metric is the relative average fees per address, compared to the average in the 50 days preceding Arbitrum’s airdrop, shown in Figure 8. One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other platforms.
Figure 8. Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
Figure 9. Total fees relative to the average for the 50 days before Arbitrum’s airdrop.
Observation 7: Arbitrum’s total daily fees spiked around the time of the airdrop. Figure 9 shows that Arbitrum’s airdrop did not confer a substantial advantage over its competitors when it comes to its income stream from transaction fees. Although the total amount of fees it earned on the day of the drop reached a peak, this was momentary.
Indeed, at the relative lead, Arbitrum held over Optimism with respect to this metric, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
Figure 10. Daily TVL increase relative to the average for the 50 days before Arbitrum’s airdrop.
4.3. Total Value Locked
The Total Value Locked (TVL) of a platform is a metric that measure the combined value of all assets stored in that platform.
[…]
5. Common Pitfalls
Airdrops share similarities with traditional loyalty programs, such as newcomer bonuses offered to new customers by banks and credit card companies. Furthermore, they also share some common pitfalls. The blockchain setting and the mechanics adopted by most airdrops may exacerbate these pitfalls, or introduce new ones.
Airdrop farmers. Sophisticated users known as airdrop farmers employ elaborate tactics to maximize the amounts of airdrop tokens they receive. Blockchain platforms have adopted various measures to curb the possibility of airdrop farmers gaming the system and obtaining an inappropriate amount of rewards. For example, a popular approach is to limit the amount of rewards that a single user can receive. Alas, a single user can pose as multiple users, in a move called a sybil attack.
Thus, platforms resort to Proof-of-Humanity services, such as Gitcoin passport. These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human. Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH (Gitcoin, 2023). Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop (Barthere, 2023), but this may lead to false negatives (X-explore, 2023). There are some recent examples of platforms like Celestia considering using sybil resistance solutions (Labs, 2023c).
There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions (Protocol, 2023), to sharing posts on social media (Labs, 2023b). These tasks sometimes appear arbitrary, leading to frustration among users.
The problem with the aforementioned methods is that they are easy to automate (Jumper-sketch, 2023) and cheaply fooled, even more-so when platforms offer retroactive rebates on transaction fees (Optimism, 2023). Furthermore, the reliance of many platforms on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops. Even using biometric identification does not guarantee sybil-proofness (TrustaLabs, 2023). A different approach used by platforms is to announce airdrops that retroactively reward users that were active before the announcement. But, farmers can prepare in advance by interacting with these platforms even when no airdrop is officially announced (Copeland, 2023), as was the case in dYdX’s airdrop (Edwards, 2022).
The phenomena of reward farming is not limited to crypto-related airdrops, and similar ones can be found in “traditional” loyalty programs. In particular, the practice of credit card churning, where users apply for credit cards for the sole purpose of receiving the rewards bestowed upon newcomers, only to cancel the cards once the rewards were received. Given that similar problems exist even in traditional contexts, where users can be easily identified and penalized, it seems that reward farming cannot be solved easily.
Distributing governance tokens may be dangerous. Some platforms airdrop governance tokens, as a means to decentralize their governance process. Such tokens allow holders to vote in the platforms’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value (Dotan et al., 2023).
There is some empirical evidence that suggests that airdrops of governance tokens may be more successful than airdrops of non-governance tokens. In a recent analysis of the market capitalization of crypto tokens performed by Makridis et al. (Makridis et al., 2023), the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped. But, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system (Messias et al., 2023a; Feichtinger et al., 2023; Sharma et al., 2023). Also, to users that do not have the platform’s best interests in mind, and may even act to the determent of the platform.
6. Design Guidelines
The pitfalls outlined earlier, while indeed sobering, may inform platforms embarking on the airdrop journey, and can shed light on the path to potential success.
Alternative incentives may better sustain user engagement. An airdrop’s long-term benefit to a platform can be tricky to measure, given that the potential benefits may be indirect and hard to quantify in advance, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from platforms to users and is thereafter no longer available for other uses. Furthermore, some costs and effects cannot be easily foreseen, like those arising from distributing governance tokens to users.
Instead of using airdrops, a community could adopt other measures which have a more predictable relation between costs and benefits. For example, a very natural alternative to airdrops is for the community to vote to programmatically reward power users with discounts on future interactions. When considering L2s, these discounts can be given on transaction fees. Thus, if a user receives this reward, it necessarily has to use the platform again in the future to benefit from the incentive, thus effectively fostering continued user engagement. Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the platform, and the platform’s costs go solely towards users actively using the system.
Nevertheless, while fee discounts are more focused towards incentivizing continued interaction with the platform, they are not without their own challenges. For example, one should carefully design a proper discount mechanism, covering both the eligibility criteria for users, and the height of awarded discounts. Furthermore, it is not immediately apparent that discounts can attract users to a similar degree as the more immediate and tangible rewards offered by airdrops. Another option is to perform multiple airdrops over a relatively long period of time, instead of a single one.
Target well-known, reputable entities. Instead of rewarding pseudonymous users, one can target DAOs that are building relevant apps. In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs (Foundation, 2023a). Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students (Foundation, 2023c) or members of the technical community (Foundation, 2023b) who wish to take part in research and build tools relevant to the platform. Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world (Foundation, 2021).
Proactive oversight and community engagement. While the airdrop is ongoing, the platform’s data should be carefully analyzed to verify that no one is attempting to game the system. A good example was put forth by Linea, whose team uncovered a vulnerability (Liena, 2023) that allowed users to cheat their incentive mechanism. The prompt discovery prevented cheaters from obtaining over one-third [^2] A dashboard is available on Dune at https://dune.com/rusty378/guildxyz-pins. of non fungible tokens (NFTs)offered as incentives.
A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities (Team, 2023), whether these are actively used or not. For example, a disclosure by an AzukiDAO community member allowed the platform to mitigate a vulnerability relatively quickly (MetaSleuth, 2023). A platform’s oversight should extend beyond on-chain data. In particular, social media is rife with scammers who advertise “fake” airdrops for their platforms, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds. Even platforms that are not planning an airdrop are targeted by such ploys.
Actively engaging with the community regarding technical matters can prove beneficial, as well. For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar (cygaar, 2023), who found that it could be implemented more cost-effectively.
Starkware’s governance team have so far performed admirably in this regard (Chevet, 2023). The team organized “town hall meeting”-esque discussions and workshops, and actively invited community members to come, participate and voice their opinions. These events provide an effective and transparent communication channel, directly reaching active community members. Operating transparently and giving the community visibility into the inner-workings of the platform can assure that even if technical issues arise, they are met with understanding by users.
Rewards should scale with costs. The ramifications of Goodhart’s law are readily observable in previous airdrops. For example, some airdrops attempt to incentive user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the platform in the past. Consequently, users can game the system by performing meaningless wash-trading transactions, meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the platform.
This is further compounded by the fact that action-based metrics used by airdrops do not usually account for the actual costs incurred by users for each action. For instance, if the number of transactions is used as a metric, one should consider that if transaction fees are low, an airdrop farmer may cheaply perform the required amount of transactions. On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop. To remedy these issues, rewards should scale with the costs that are incurred by users.
7. Related Work
The majority of recent papers on airdrops concern themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
Studies. Makridis et. al. (Makridis et al., 2023) investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges. They find that airdrops of governance tokens, in particular, are positively associated with growth in market capitalization and volume.
Lommers et. al. (Lommers et al., 2023) offer a broad treatment of the different kinds of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based). They offer examples of projects like Blur and the pitfalls of their airdrops, together with some advice on airdrop implementation strategies. In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success and we provide practical guidance on how to toggle these.
Fan et. al.(Fan et al., 2023) study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristic of community members and the effectiveness of airdrop. They find that users receive more rewards tend to take positive actions towards the community. They summarize several arbitrage patterns and confirm the current detection is not sufficient in screening out airdrop hunters. They use an unsupervised hierarchical clustering method to capture the behaviors of eligible airdrop addresses after receiving rewards and to divide them into interpretable categories.
Allen (Allen, 2023) examines airdrop design as co-evolutionary process and makes three contributions. First, applying a co-evolutionary lens to airdrops, where participants learn and adapt their strategies (e.g., new criteria, on-chain farming) both within and between web3 ecosystems. Second, nine airdrop cases studies (e.g., Optimism, Arbitrum, Blur) that demonstrate the expanding set of airdrop strategies (e.g., task-based claiming). Third, implications for airdrop design including more feedback loops to facilitate learning (e.g., multi-round airdrops), ongoing innovation in airdrop criteria (e.g., randomness), and a reversion back to simple designs for some projects (because of the costs of airdrop design).
Allen et. al.(Allen et al., 2023) consider why projects might choose this unusual and costly means of token distribution. It considers a selection of high-profile airdrops as case studies between 2014 and 2022. This is the first comprehensive analysis of the rationales and mechanisms of Web3 token airdrops. We find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralisation of ownership and control of a project (building community, providing regulatory protection, and enhancing security). While many commentators describe airdrops as primarily for marketing, they claim that marketing is a weak rationale for airdropping tokens and the evidence of successful marketing-driven airdrops lacking.
Technical Aspects of Airdrops. Frowis et. al. (Fröwis and Böhme, 2019) point out that efficient transfers to a large number of recipients in the case of airdrops present issues on Ethereum. They find that cost savings of a factor of two are in fact possible, but require specific provisions in the smart contract implementing the token system. Pull-based approaches, which use on-chain interaction with the recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient. Total costs are broadly linear in the number of recipients independent of the technical approach.
Wahby et. al. (Wahby et al., 2020) point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities. In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely-used user credentials (ECDSA and RSA). Private airdrop for RSA builds upon a new zero-knowledge argument of knowledge of the factorization of a committed secret integer, which may be of independent interest. Their fastest implementation takes 40–180 ms to generate and 3.7–10 ms to verify an RSA private airdrop signature.
8. Conclusions
In this work we conducted analysis for of the common pitfalls of airdrops and propose concrete ways to improve the situation. Specifically, we analyze the data pertaining to four recent airdrops and conclude that the majority of proceeds (as high as 95%, typically only within a couple of transfer steps) have been sold via exchanges, implying that the airdrops fall far short of their intended purpose. We hope that the suggestions in this paper provide inspiration for researchers and practitioners alike.
Appendix A Reproducibility
To ensure complete reproducibility of our empirical analysis, we have openly shared both our dataset and scripts (Messias et al., 2023b). The primary data source employed in Section 3 comprises blockchain data extracted from both Ethereum and zkSync Era archive nodes. We meticulously collected all blocks, transactions, and events associated with the contracts of the four airdrops under study.
For Section 4, we sourced data from Growthepie.xyz (Growthepie, 2023). This platform provided various metrics (daily transaction count, daily active addresses, median transaction cost, total value locked, fees paid by users, stablecoin market cap) for several key platforms (zkSync Era, OP Mainnet, Arbitrum, Immutable X, Polygon zkEVM). All these metrics were smoothed using a 7 day moving average.
3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@
-Airdrops: Giving Money Away Is Harder Than It Seems
-Abstract.
-Airdrops are used by blockchain applications and platforms to attract an initial user base, and to grow the user base over time.
-In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying platform, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop has been completed.
+
+Abstract
+Airdrops are used by blockchain applications and protocols to attract an initial user base, and to grow the user base over time.
+In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying protocol, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop.
 Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
@@
 Our analysis is followed by an overview of common pitfalls that common airdrop designs lend themselves to, which are then used to suggest concrete guidelines for better airdrops.
-1. Introduction
-Blockchain platforms commonly create reward schemes to lure new customers and maintain the loyalty of existing users.
-In recent years, giving platform-minted tokens to users, or airdropping tokens, has proven itself as a particularly popular scheme (Makridis et al., 2023).
-Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives.
-Intuitively, this is not optimal, and may result in wasting funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
-While the basic premise of airdrops is relatively straightforward, the design space for such reward schemes is large, and may depend on the characteristics of the platform designing the airdrop.
-For example, some airdrop mechanisms target power-users and endow them with a generous amount of rewards in the hopes that these users will generate valuable economic activity, and thus attract other users as well.
+keywords:
+Airdrop design, Airdrop farmer, Web3, Layer 2
+1 Introduction
+Blockchain protocols often design reward programs to attract new customers and foster the loyalty of existing users.
+In recent years, the practice of distributing platform-minted tokens, known as “airdropping,” has gained widespread popularity [53].
+For example, in 2023 alone, 4.56 billion US dollars worth of tokens were airdropped to users across various protocols [9].
+Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing that there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives.
+Intuitively, this is not optimal and may result in the loss of funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
+Although the fundamental concept of airdrops is relatively straightforward, the design space for such reward schemes is expansive and may vary depending on the characteristics of the platform implementing the airdrop.
+For instance, certain airdrop mechanisms focus on power users, granting them a substantial amount of rewards with the expectation that these users will stimulate valuable economic activity, subsequently attracting additional users.
+However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance.
+It allows users to suggest or vote on changes within a platform, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token.
+This approach introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making voting power [56, 22, 68].
 To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops.
 Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process.
-Specifically, we analyze the data pertaining to four recent airdrops and conclude that the majority of proceeds (as high as 95%, typically only within a couple of transfer steps) have been sold via exchanges, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers as opposed to regular users.
-This comparison is followed by a description of common pitfalls faced by airdrops in the past.
-Given that the blockchain ecosystem is still quite young, it is to be expected that the theory and practice of blockchain airdrops is still in its infancy, and thus that previous airdrops may have not been entirely successful.
-Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms which are more fair to honest users.
-1.1. Contributions
-Arbitrum study.
-We perform a comprehensive case study of the Arbitrum (Kalodner et al., 2018) airdrop by measuring elements such as transaction volume, token, distribution structure, and token value before and after the airdrop period.
-Quantitative analysis.
-We quantitatively analyze airdrops executed by ENS (ENS, 2023), dYdX (Juliano, 2018), Gemstone (Haig, 2023; lingland09, 2023), and 1inch (Bukov and Melnik, 2020).
-We show that most of the funds received through these airdrops are sold on exchanges, instead of being used to create dapp or platform engagement on the part of end-users.
-Qualitative analysis. We qualitatively analyze previous airdrops, and propose guidelines for developing future airdrops in an effort to avoid some of the pitfalls in the airdrops we analyze.
-Scientific reproducibility.
-To ensure the scientific reproducibility of our empirical analysis, we share our dataset and scripts (Messias et al., 2023b).
-Refer Section A in the appendix for more details.
-2. Airdrop Desiderata
-Airdrops are typically used to bootstrap a community of users.
-We break down this overarching high-level goal into multiple sub-goals.
+Specifically, we analyze the data pertaining to five popular airdrops (ENS [19], dYdX [39], 1inch [10], Arbitrum [41], and Uniswap [1]) and a fake airdrop conducted by a sybil farmer (Gemstone [35, 50]).
+We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of airdropped tokens that they receive.
+Furthermore, we describe the common challenges faced by airdrops in the past.
+Given that the blockchain ecosystem is still quite young, it is expected that the theory and practice of blockchain airdrops are still in its infancy and thus that previous airdrops may not have been entirely successful in the long run.
+Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms that are more fair to honest users.
+1.1 Research Questions
+In this section, we address three critical research questions central to our study.
+[…]
+We evaluate the effectiveness of blockchain protocols in converting short-term user incentives, such as airdrops, into long-term user engagement.
+The primary goal of conducting an airdrop is to attract users in the short-term and encourage them to become long-term users.
+Our investigation aims to measure how well these protocols achieve this objective.
+Contrary to the assumption that airdrops are the primary driver of long-term user base growth, our research reveals otherwise.
+Through an exploration of past airdrops conducted, we demonstrate that airdrops may not be the sole determinant of sustained user retention.
+For example, findings from a recent report by Messari [8] indicate that users join a platform when short-term incentives are available, such as airdrops.
+However, these protocols struggle to retain users in the long-term, as users often shift their attention to protocols offering better incentives as soon as the initial rewards expire.
+This behavior implies that protocols successfully onboard users during incentive periods, like airdrops, but witness a drop in user engagement once the incentives come to an end.
+Additionally, protocols not conducting airdrops also manage to onboard users due to an expectation of potential future airdrops.
+Our findings align with this observation, as our analysis indicates that, even after Arbitrum conducted an airdrop, other protocols attracted user attention and increased their total value locked (TVL).
+This suggests that airdrops alone may not be sufficient to ensure sustained user retention of the platform performing the airdrop, and other factors play crucial roles in long-term platform success.
+[…]
+Since protocols face challenges in achieving sustained user retention through airdrops alone, an alternative strategy involves incentivizing developers and innovators to build new applications, such as decentralized exchanges (DEXs)and non-fungible token (NFT).
+This approach aims to foster a dynamic ecosystem that benefits all users on the platform.
+To enhance user experience and maintain user loyalty, a focus on developing more, potentially exclusive, decentralized applications (dApps) is crucial.
+Another interesting approach is to provide additional incentives, such as discounts on fees for future platform interactions.
+For example, in the context of a rollup, users who receive airdropped tokens could redeem them for discounted fees, encouraging long-term engagement.
+Also, implementing multi-round airdrops allows protocols to learn from previous rounds, enabling iterative improvements to address previously identified issues [2].
+Community engagement also plays a vital role in understanding the needs and preferences of users.
+In line with these considerations, Blast has introduced an innovative airdrop strategy centered around user contributions to their protocol [55].
+This strategy involves awarding points to users as a quantitative measure of their contribution.
+Points are granted for actions such as successfully bringing in new participants to the blockchain protocol, similar to a referral system, or for users who bridge (send) their tokens from one chain to Blast.
+Furthermore, user engagement in blockchain protocols often hinges on short-term incentives.
+Users tend to join the platform when these incentives, such as airdrops, are available.
+However, a notable trend is the propensity of users to migrate opportunistically from one blockchain to another, following the trail of incentives.
+Consequently, a platform experiences an influx of users during incentive periods, only to witness a decline when these incentives cease.
+[…]
+A crucial aspect to consider involves airdrop farmers.
+These are sophisticated users who employ intricate tactics to maximize their chances of receiving more tokens per airdrop.
+They often create numerous wallet addresses controlled by a single entity, a practice known as Sybil [37, 64].
+Detecting Sybils can be challenging [21, 51, 36, 73, 59], especially when these accounts execute automatic and randomized actions (e.g., token transfers, bridging, providing liquidity, exchanging tokens, NFTs, etc.) on the blockchain without linking their wallets.
+Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such actions [70, 29, 12, 15].
+Given its intricacy, current methods for identifying Sybils predominantly rely on graph network analysis [21, 51] and machine learning techniques [36, 73, 59].
+Another common challenge involves airdropping governance tokens.
+These tokens empower holders to propose or vote on protocol changes.
+Distributing these tokens can be risky when users do not prioritize the protocol’s best interests.
+Additionally, it poses the danger of granting excessive power to certain users, as these protocols typically use a one-token-one-vote mechanism [56, 22, 68].
+While acknowledging the challenges highlighted earlier, we believe that these challenges can serve as valuable lessons for protocols venturing into the realm of airdrops, offering insights into the potential path to success.
+Furthermore, it is essential to uphold a positive reputation within the community by avoiding any occurrences of insider trading [5, 42].
+Insider trading refers to individuals within an organization using privileged information to gain a financial edge over other protocol users.
+This practice is a common violation of securities laws in traditional financial markets [6], and engaging in it can lead to negative reactions from the blockchain community [69, 61, 23].
+Finally, the points discussed above raise questions about whether protocols are effectively achieving their goals post their airdrop campaigns.
+1.2 Summary of our Contributions
+We summarize our contributions as follows.
+[…]
+[…]
+[…]
+[…]
+2 Airdrop Desiderata
+Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with the blockchain protocol or application.
+They are widely utilized for these purposes, with numerous examples documented in the literature (see Table 1).
+Any protocol capable of creating tokens can utilize airdrops to distribute them to users.
+For example, blockchain rollup solutions like Arbitrum and Optimism, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
+Airdrops can manifest in various forms, with single-round and multi-round distributions being the most prevalent [2].
+In a single-round airdrop, tokens are distributed to users all at once, whereas multi-round airdrops involve distributing tokens over multiple rounds each following different strategies.
+This approach allows leveraging insights from previous rounds to address challenges encountered, such as mitigating potential sybil (i.e., multiple accounts controlled by a single entity) by observing past user behavior patterns [2].
+The choice between single-round and multi-round airdrops depends on the protocol’s objectives and community dynamics.
+Community engagement is a critical aspect of airdrop development; it should be carefully considered by protocol developers.
+Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens.
+Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner.
+This difference in user base size introduces an added layer of complexity when performing sybil detection, as there are more accounts to evaluate and potentially safeguard against.
+If not performed correctly, this could impact community happiness.
+In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community.
+However, addressing industrial farming poses significant challenges and complexities.
+To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils.
+Under this system, sybils have the option to self-report and receive 15% of their designated token allocation.
+Failure to self-report may result in them not receiving tokens if they are identified by the company or its affiliates [46].
+Table 1: List of popular airdrops we analyze in this paper.
+[…]
+[…]
+dYdX Sept. 2021 Still active Ethereum Continuous DEX
+1inch Dec. 2020 Still active Ethereum Continuous DEX
+[…]
+Arbitrum Mar. 2023 Sept. 2023 Arbitrum One Single Single
+Uniswap Sept. 2020 Still active Ethereum Continuous DEX
+[…]
+Next, we break down airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals.
 These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
-1.
 Attract users in the short-term.
-Historically, airdrops were used by new blockchain platforms to create an initial user base.
-In particular, decentralized platforms commonly become more useful and attractive to users as a function of the amount of economic activity taking place.
-2.
+Historically, airdrops have been employed by new blockchain protocols to establish an initial user base and to provide an initial boost of liquidity to the underlying chain and its protocols.
+Decentralized platforms, in particular, tend to become more useful and attractive to users as the level of economic activity increases.
 Foster long-term involvement.
 Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time.
-Ideally, platform users should become regular, everyday users.
-This can be achieved by issuing rewards that can only be used within the platform, similarly to frequent-flyer points.
-For example, in layer 2 blockchains (L2s) (Sguanci et al., 2021), one may award discounts on future transactions.
-Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the platform has to offer.
-For example, Linea’s airdrop quests gave users an in-depth tour of the platform’s various features and use-cases (Labs, 2023a).
-3.
+Ideally, users should become regular, everyday users.
+This can be achieved by issuing rewards that can only be used within the blockchain protocol or application, similarly to frequent-flyer points.
+For example, in layer-2 blockchains (L2s) [67], one may award discounts on future transactions.
+Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the blockchain has to offer.
+For example, Linea’s airdrop quests gave users an in-depth tour of their various features and use-cases [44].
 Target value-creating users.
-An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of the platform.
-In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools (Yaish et al., 2022) and decentralized exchanges (DEXs) (Yaish et al., 2023), or across a variety of tokens.
+An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform.
+In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [78] and DEXs [77], or across a variety of tokens.
 In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable.
-Such users provide additional use-cases for a platform, and thus serve to attract other users.
-Table 1. Airdrops we analyze in this paper.
-Airdrop Start date End date Blockchain
-ENS October 2021 May 2022 Ethereum
-dYdX September 2021 Still active Ethereum
-1inch December 2020 Still active Ethereum
-[…]
-3. Exchange Aftermarket Analysis
+Such users provide additional use cases for a platform, and thus serve to attract other users.
+3 Exchange Aftermarket Analysis
 The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place.
 Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them.
-Thus, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them (jhackworth, 2022), and 61% have sold their tokens shortly after the drop in the case of ParaSwap (Fan et al., 2023).
-In both cases, recipients mostly stopped using the corresponding platforms within months.
-This behavior indicates that the airdrop did not succeed in retaining recipients, or that recipients are mostly sybil accounts.
-Furthermore, a quick sell-off may alarm the market, if it is interpreted as a sign of a lack of belief in the long-term success of the platform.
-Below we consider data related to the four airdrops (ENS, dYdX, Gemstone, and 1inch) in Table 1 that were collected from an Ethereum and zkSync Era archive nodes.
-We have made available both our data collection and analysis scripts, along with the dataset itself (Messias et al., 2023b).
-To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query:
-select * from addresses_ethereum.dex
-Platform Token Tokens % sold at # of recipient Tokens received per address
-supply distributed exchanges addresses Mean Std. Median Min Max
-ENS $1\times 10^{8}$ $1.96\times 10^{7}$ 36.51% $102822$ $190.93$ $141.57$ $152.4$ $0.0001$ $1143.54$
-dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.35% $42682$ $2325.85$ $91336.75$ $392.56$ 00 $2.47\times 10^{7}$
-1inch $1.5\times 10^{9}$ $8.3\times 10^{7}$ 53.47% $43402$ $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
-Gemstone $5.06\times 10^{16}$ $5.04\times 10^{16}$ 95.01% $20014$ $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
-Table 2. Airdrop distribution statistics for the four airdrops.
-3.1. Tokens Distributed Over Time
-Table 2 proves some numbers on how airdropped tokens are spread across the user base.
-We observe that the number of recipients is quite high for every analyzed airdrops, which may indicates some effort on the part of airdrop farmers (column 3).
-Moreover, the data indicates that the tokens were often traded on exchanges, as evident from the fact that the first transfer made by the addresses after the airdrop was a sale in an exchange (column 10).
-Gemstone stands out from the rest, with 95% of its tokens ending up on exchanges.
+For example, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them [38], and 61% have sold their tokens shortly after the drop in the case of ParaSwap [21].
+In both instances, the majority of recipients ceased using the relevant blockchain protocols within a few months.
+This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely sybil accounts.
+Additionally, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects.
+Here, we examine data associated with six airdrops (ENS, dYdX, 1inch, Arbitrum, and Uniswap, including Gemstone) presented in Table 1, which were gathered from archive nodes of Ethereum and zkSync Era.To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query: select * from addresses_ethereum.dex and from Etherscan [20].
+Table 2: General airdrop distribution statistics.
+[…]
+supply distributed exchanges addresses
+[…]
+dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$
+1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$
+Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$
+[…]
+Table 3: Tokens received per address.
+[…]
+[…]
+dYdX $2325.85$ $91336.75$ $392.56$ 00 $24.7\times 10^{6}$
+1inch $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
+Gemstone $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
+Arbitrum $1874.02$ $1566.92$ $1250$ $625$ $10250$
+Uniswap $620.13$ $9347.91$ $400$ $400$ $2.1\times 10^{6}$
+[…]
+3.1 Tokens Distributed Over Time
+Table 2 presents the number of airdrops distributed and subsequently sold on exchanges.
+Additionally, Table 3 details the distribution of airdrops per user.
+These tables provide quantitative insights into how airdropped tokens are disseminated across the user base.
+Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2).
+Furthermore, the data indicate that tokens were frequently traded on exchanges.
+This is supported by the observation that the initial transfer made by the addresses post-airdrop was often a sale on an exchange (see column 4 in Table 2).
+Notably, the case of Gemstone is particularly striking, with 95% of its tokens being sold on exchanges.
+In this particular Gemstone case, a non-open sourced DEX created by the airdrop farmer [35].
 Furthermore, Gemstone’s airdrop stands out significantly in terms of the number of tokens distributed compared to other airdrops.
-This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops.
+This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops as shown in Table 3.
 Remarkably, Gemstone allocated the lion’s share, specifically 99.53%, of their total token supply during their airdrop campaign.
 It is important to highlight that Gemstone was executed primarily as a sybil attack rather than a legitimate airdrop.
-Platform Airdrop recipients # of hops
-Total # sent to exchanges % sent to exchanges Mean Std. Median Min. Max.
-ENS $102822$ $86157$ 83.79% 1.75 1.06 2 1 50
-dYdX $42682$ $41320$ 96.81% 2.25 0.59 2 1 18
-1inch $43402$ $41913$ 96.57% 1.85 0.64 2 1 7
-Gemstone $20014$ $20000$ 99.93% 1 1 1 1 1
-Table 3.
-Airdrop recipients’ interaction with exchanges.
-Gemstone stands out as an exception with a median of 1 hop, while other protocols show a median of 2 hops for connecting airdrop recipients with exchanges.
+Figure 1: Tokens claimed per day for each of the Airdrops analyzed.
+Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users.
+Figure 1 illustrates the distribution of accounts claiming their tokens each day.
+Arbitrum conducted a substantial airdrop, distributing $1162166000$ ARB tokens among $625143$ selected accounts.
+Notably, $583137$ (93.28%) of these accounts successfully claimed 94.03% of the allocated ARB tokens.
+Remarkably, 72.45% of the accounts completed their token claims on the first day, with an additional 14.41% doing so on the following day.
+This cumulative figure amounts to almost 87% of accounts claiming the Arbitrum airdrop within one day of its release, indicating high user engagement and prompt action among the majority of participants.
+Table 4: Airdrop recipients’ interaction with exchanges.
+[…]
+Total # sent to exchanges % sent to exchanges
+[…]
+dYdX $42682$ $41320$ 96.81%
+1inch $43402$ $41913$ 96.57%
+Gemstone $20014$ $20000$ 99.93%
+Arbitrum $583137$ $555040$ 95.18%
+[…]
+Table 5: Distribution of hops for airdrop recipients connecting with exchanges.
+[…]
+Mean Std. Median Min. Max.
+[…]
+dYdX 2.25 0.59 2 1 18
+1inch 1.85 0.64 2 1 7
+Gemstone 1 1 1 1 1
+Arbitrum 2.71 $0.79$ $3$ $1$ $32$
+Uniswap 2.11 $0.43$ 2 1 9
+[…]
 (a) Over the number of post-airdrop blocks.
 (b) Over the number of post-airdrop days.
-Figure 1. Time for airdrop tokens to be transferred to exchanges.
-Focusing on Table 3, we computed the shortest path taken from each recipient address of an airdrop to any of the exchange address in our set.
-Our findings reveal that the transfer of tokens to exchanges typically takes only a few steps, suggesting that airdrop farmers, in general, are not making significant efforts to hide their activities.
-The outlier is Gemstone where in all cases tokens were sold within only one hop.
-Surprisingly, the majority of accounts managed to reach an exchange with a relatively low number of intermediating hops.
-This observation underscores the prominent role that exchanges play within the crypto ecosystem.
+Figure 2: Time for airdrop tokens to be transferred to exchanges.
+Users typically interact with exchanges to convert one token for another and to sell them.
+To assess how often users receiving an airdrop sell these tokens on exchanges for profit, we analyzed user interactions with exchanges post-airdrop.
+Table 4 indicates that a significant proportion of airdrop recipients engage with exchanges to trade their tokens.
+This engagement ranges from a minimum of 83.79% in the case of ENS to a maximum of 99.93% for Gemstone.
+Additionally, Table 5 presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset.
+Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities.
+Gemstone is a notable outlier, with all tokens being sold within a single hop.
+Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (between 2 and 3 hops).
+This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
 (a) ENS.
@@
 (d) 1inch.
-Figure 2. Network structure for selling airdropped tokens in 4 airdrops.
-Figure 1 shows the CDFs that demonstrate how quickly tokens are sold at changes.
-Per Figure 0(a), we see that within about a million blocks, the majority of tokens are gone.
-For Gemstone the number of blocks significantly higher than others because of a delay introduced by airdrop farmers.
-As the block generation is quite different for Ethereum (a new block is created every 15s on average) and zkSync, we normalized the blocks in terms of the number of days.
-Per Figure 0(b), within a single day, 64.67% of the 1inch tokens were sent to an exchange.
-In comparison, ENS had a transfer rate of 45.86%, while dYdX stood at 37.41%.
-This does not align with one of the primary objectives of a protocol when conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2.
-3.2. Token Distribution Graphs
+(e) Arbitrum.
+(f) Uniswap.
+Figure 3: Network structure for selling airdropped tokens in 6 airdrops.
+Figure 2 shows the cumulative distribution functions (CDFs) demonstrating how quickly users interact with an exchange right after receiving tokens from an airdrop.
+As shown in Figure 2(a), the majority of accounts exchanged their tokens within approximately one million blocks.
+The number of blocks for Gemstone is significantly higher than others due to a delay introduced by airdrop farmers.
+Given the difference in block generation times between Ethereum (with a new block created approximately every 15 seconds) and zkSync, we normalized the blocks in terms of days.
+According to Figure 2(b), within a single day, 64.67% of the accounts interacted with an exchange.
+In comparison, ENS had an interaction rate of 45.86%, while dYdX stood at 37.41%.
+This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2.
+The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
+3.2 Token Distribution Graphs
 To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network.
@@
 Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges.
-To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdropped the first few hours after the airdrop and plot their largest component in Figure 2.
-We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer.
-We show that for all protocols but Gemstone, Uniswap is the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from the addresses receiving the airdrops, followed by SushiSwap.
+Lastly, Arbitrum accounts for $921650$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
+To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Figure 3.
+We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [20].
+We show that for all protocols but Gemstone, the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.
 […]
-(a) Absolute number of unique daily active addresses.
-(b) Daily active addresses, relative to the average for the 50 days before Arbitrum’s airdrop.
-Figure 3. Unique daily active addresses per platform. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
-Figure 4. Median daily transaction fees, in USD.
-4. Measuring the Airdrop Lift
-Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially.
-Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic.
+(a) Absolute number.
+(b) Relative daily active addresses.
+Figure 4: Unique daily active addresses per protocol: (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
+Figure 5: Ratio between the number of unique daily active addresses in Arbitrum and Optimism.
+(a) Daily median fees.
+(b) Daily average fees.
+(c) Average fees spent per unique address per day.
+Figure 6: Daily transaction fees in USD. (a) Median; (b) Average; and (c) Per unique address.
+4 Measuring the Airdrop Lift
+Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [21].
+Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [30, 21].
 In this section, we examine this through the lens of some relevant metrics applied to the Arbitrum airdrop.
-We rely on data source from Growthepie (Growthepie, 2023) and list details in Section A in appendix.
-4.1. Unique Active Addresses
-Figure 3 examines the number of unique addresses that were active on each platform using two different approaches.
+We rely on data source from Growthepie [34].
+This platform supplies various metrics, including daily transaction count, daily active addresses, median transaction cost, TVL, fees paid by users, and stablecoin market cap, for several key protocol such as zkSync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM.
+To enhance clarity, all these metrics were subjected to a 7-day moving average smoothing process.
+4.1 Unique Active Addresses
+Figure 4 examines the number of unique addresses that were active on each protocol using two different approaches.
 The Figure 4(a) depicts the absolute number of addresses.
-The Figure 2(b) shows the relative amount of unique daily active addresses for each platform, when compared to the same platform’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 (Rustgi, [n.
-d.]).
-This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users (Barthere, 2023).
-Observation 1: others outperformed Arbitrum, without an airdrop.
+The Figure 4(b) shows the relative amount of unique daily active addresses for each protocol, when compared to the same protocol’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 [66].
+This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users [7].
+{observation}
+others outperformed Arbitrum, without an airdrop.
 The absolute number of addresses has increased for Arbitrum post-airdrop, and remained at least 50% higher than its pre-airdrop level in the four months since then.
-But, by looking at other platforms, it becomes apparent that a similar result can be attained without an airdrop.
-This is demonstrated by Optimism’s trajectory in Figure 2(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 (Nijkerk, 2023).
+However, by looking at other protocols, it becomes apparent that a similar result can be attained without an airdrop.
+This is demonstrated by Optimism’s trajectory in Figure 4(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 [62].
 This is further exemplified by zkSync Era overtaking Arbitrum’s in Figure 4(a), two months after the airdrop.
-Observation 2: fees may explain the closing bap between Arbitrum and Optimism.
+{observation}
+fees may explain the closing gap between Arbitrum and Optimism.
 Arbitrum leads over Optimism with respect to the number of unique daily active addresses.
+We turn our focus to Figure 5, which depicts Arbitrum’s lead over Optimism with respect to the number of unique daily active addresses.
 While in the 50 days before the airdrop Arbitrum has enjoyed an average of 2.6 times more unique daily active addresses than Optimism, this gap has recently started to close: the average over the last 50 days in our dataset is considerably lower, and is equal to 1.83 times more.
-Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 4.
-Observation 3: the number of unique addresses oscillates over time.
+Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 6(a).
+We note that with respect to the average fee, the gap between the protocols is even smaller over the same period as shown in Figure 6(b).
+{observation}
+the number of unique addresses oscillates over time.
 Some of the previous observations exhibit oscillatory behavior, in particular in Figure 4(b) depicting the relative amount of unique addresses: rising, reaching a peak, and then subsequently dropping, all in rapid succession.
 In particular, the phases of Optimism and Arbitrum are inverted regarding the relative amount of unique addresses.
-One may attempt to explain this by reasoning that when a peak is reached in one of the platforms, fees should increase due to the load, thus prompting cost-conscious users to turn to a different platform.
-But, when considering the median transaction fees paid per day (per Figure 4), it is apparent that fees do not have corresponding peaks and dips.
-In fact, Arbitrum consistently had lower median fees mid-May.
-Caveat: the unique address metric is gameable.
-The overall picture painted so far is that with regard to the daily unique account metric, Arbitrum’s airdrop did not succeed in generating long-lasting user involvement.
-An important caveat of this metric is that in blockchains, users can typically costlessly create and use multiple addresses.
-This practice is also called a sybil attack, with attackers also known as sybils.
-Sybil attacks may benefit users, as some airdrops limit the amount of tokens that a single address can receive.
-Thus, using the number of unique addresses as a metric may paint an incomplete picture: the large changes we observed may be caused by sybil attackers.
-Although some employ sophisticated tools to identify sybil addresses, these are heuristic in nature and may fail to capture all of them.
-Thus, we need to consider other metrics that are hopefully more sybil-resistant and representative of genuine user activity.
-4.2. Transaction-related Metrics
+One may attempt to explain this by reasoning that when a peak is reached in one of the protocols, fees should increase due to the load, thus prompting cost-conscious users to turn to a different protocols.
+But, when considering the median transaction fees paid per day (per Figure 6(a)), it is apparent that fees do not have corresponding peaks and dips.
+In fact, Arbitrum consistently had lower median fees in mid-May.
+[…]
+4.2 Transaction-related Metrics
 A natural next step is to examine transaction-related metrics.
-In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee (Gafni and Yaish, 2022), when ignoring retroactive rebates or the possibility of platforms operating at a loss (Optimism, 2023).
-Observation 4: the Arbitrum-Optimism gap is smaller, when considering transactions.
-Figure 4(a) depicts the absolute daily transaction count per platform, and Figure 4(b) shows the relative amount of transaction count for each platform compared to the same platform’s average transaction count for the period between February 1st and March 22nd, 2023.
-These figures are quite similar to the corresponding ones for the unique address count metric.
+In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee [32], when ignoring retroactive rebates or the possibility of blockchain protocols operating at a loss [63].
+{observation}
+[…]
 Some notable differences are that with respect to transaction counts, the gap between Arbitrum and Optimism has nearly closed entirely by the end of July.
 Furthermore, Figure 7(b) shows that the daily transaction count for Immutable X has almost halved since Arbitrum’s airdrop.
-Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the platform.
+Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the protocol.
 (a) Absolute daily transactions.
-(b) Daily transactions relative to the average for the 50 days before Arbitrum’s airdrop.
-Figure 5. Daily number of transactions per platform. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
-Observation 5: Arbitrum’s per address transaction count dropped since the airdrop.
-To better understand how Arbitrum’s airdrop affected user engagement in other platforms, we now present the Figure 6, which depicts the relative average daily transaction count per unique address.
+(b) Relative daily transactions.
+Figure 7: Daily number of transactions per protocol. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
+{observation}
+Arbitrum’s per address transaction count dropped since the airdrop.
+To better understand how Arbitrum’s airdrop affected user engagement in other protocols, we now present the Figure 8.
+Figure 8(b), which depicts the relative average daily transaction count per unique address.
 The figure shows that since the drop, Arbitrum’s per-user transaction count has fallen to less than 75% of its pre-airdrop level.
-Figure 6. Daily transactions per address, relative to the average before Arbitrum’s airdrop.
-Caveat: transaction counts are inaccurate when ignoring fees.
-Relying on the previous metrics may be inaccurate: a high transaction count does not necessarily signal high user engagement.
-In particular, some platforms require users to perform a certain number of transactions over some time frame to become eligible for their airdrops.
-Thus, if fees are low, a high number of transactions per account may be caused by users who perform economically-useless transactions, for the sole purpose of receiving the drop.
-This brings to mind Goodhart’s law (Manheim and Garrabrant, 2019), which posits that “when a measure becomes a target, it ceases to be a good measure”.
-Observation 6: since June, average transaction fees are similar for all platforms.
+(a) Average number.
+(b) Relative to average before Arbitrum’s airdrop.
+Figure 8: Daily transactions per address.
+[…]
+{observation}
+since June, average transaction fees are similar for all protocols.
 Conceptually, good metrics should measure a user’s amount of skin in the game.
-A possible proxy for that is the amount of funds the user is willing to expend for interacting with the platform.
-When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all platforms are quite similar in both regards.
-Furthermore, when comparing the average fee per transaction to the median one (per Figures 4 and 7), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
-Figure 7. Daily average transaction fee spent per transaction per day, in USD.
+A possible proxy for that is the amount of funds the user is willing to expend for interacting with the protocol.
+When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all protocols are quite similar in both regards.
+Furthermore, when comparing the average fee per transaction to the median one (per Figures 6(a) and 6(b)), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
 Another interesting metric is the relative average fees per address, compared to the average in the 50 days preceding Arbitrum’s airdrop, shown in Figure 9.
-One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other platforms.
-Figure 8. Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
-Figure 9. Total fees relative to the average for the 50 days before Arbitrum’s airdrop.
-Observation 7: Arbitrum’s total daily fees spiked around the time of the airdrop.
+One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other protocols.
+Figure 9: Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
+(a) Total daily transaction fees (USD).
+(b) Relative to average before Arbitrum’s airdrop.
+Figure 10: Total fees before and after the Arbitrum airdrop.
+Figure 11: Ratio between the daily fees spent in Arbitrum and Optimism.
+{observation}
+Arbitrum’s total daily fees spiked around the time of the airdrop.
 Figure 10 shows that Arbitrum’s airdrop did not confer a substantial advantage over its competitors when it comes to its income stream from transaction fees.
 Although the total amount of fees it earned on the day of the drop reached a peak, this was momentary.
-Indeed, at the relative lead, Arbitrum held over Optimism with respect to this metric, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
-Figure 10. Daily TVL increase relative to the average for the 50 days before Arbitrum’s airdrop.
-4.3. Total Value Locked
-The Total Value Locked (TVL) of a platform is a metric that measure the combined value of all assets stored in that platform.
+Indeed, per Figure 11, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
+(a) Daily TVL, in USD.
+(b) Relative to the average prior Arbitrum’s airdrop.
+Figure 12: Total Value Locked (TVL) growth: (a) Daily TVL in USD; and (b) relative to the average for the 50 days before Arbitrum’s airdrop.
+4.3 Total Value Locked
+The Total Value Locked (TVL) of a protocol is a metric that measure the combined value of all assets stored in that protocol.
+{observation}
 […]
-5. Common Pitfalls
+5 Common Airdrop Design Challenges
 Airdrops share similarities with traditional loyalty programs, such as newcomer bonuses offered to new customers by banks and credit card companies.
-Furthermore, they also share some common pitfalls.
+Furthermore, they also share some common design challenges.
 The blockchain setting and the mechanics adopted by most airdrops may exacerbate these pitfalls, or introduce new ones.
-Airdrop farmers.
+5.1 Airdrop Farmers
 Sophisticated users known as airdrop farmers employ elaborate tactics to maximize the amounts of airdrop tokens they receive.
-Blockchain platforms have adopted various measures to curb the possibility of airdrop farmers gaming the system and obtaining an inappropriate amount of rewards.
+Blockchain protocols have adopted various disproportionate measures to curb the possibility of airdrop farmers gaming the system.
 For example, a popular approach is to limit the amount of rewards that a single user can receive.
-Alas, a single user can pose as multiple users, in a move called a sybil attack.
-Thus, platforms resort to Proof-of-Humanity services, such as Gitcoin passport.
+Thus, protocols resort to Proof-of-Humanity services, such as Gitcoin passport [33].
 These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human.
-Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH (Gitcoin, 2023).
-Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop (Barthere, 2023), but this may lead to false negatives (X-explore, 2023).
-There are some recent examples of platforms like Celestia considering using sybil resistance solutions (Labs, 2023c).
-There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions (Protocol, 2023), to sharing posts on social media (Labs, 2023b).
+Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH [33].
+Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop [7], but this may lead to false negatives [76].
+There are some recent examples of protocols like Celestia considering using sybil resistance solutions [47].
+There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions [65], to sharing posts on social media [45].
 These tasks sometimes appear arbitrary, leading to frustration among users.
-The problem with the aforementioned methods is that they are easy to automate (Jumper-sketch, 2023) and cheaply fooled, even more-so when platforms offer retroactive rebates on transaction fees (Optimism, 2023).
-Furthermore, the reliance of many platforms on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops.
-Even using biometric identification does not guarantee sybil-proofness (TrustaLabs, 2023).
-A different approach used by platforms is to announce airdrops that retroactively reward users that were active before the announcement.
-But, farmers can prepare in advance by interacting with these platforms even when no airdrop is officially announced (Copeland, 2023), as was the case in dYdX’s airdrop (Edwards, 2022).
+The problem with the aforementioned methods is that they are easy to automate [40] and cheaply fooled, even more-so when protocols offer retroactive rebates on transaction fees [63].
+Furthermore, the reliance of many protocols on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops, even with biometric identification, full sybil resistance is not guaranteed [72].
+A different approach used by protocols is to announce airdrops that retroactively reward users that were active before the announcement.
+Nevertheless, farmers can prepare in advance by interacting with these protocols even when no airdrop has officially been announced [14], as was the case in dYdX’s airdrop [18].
 The phenomena of reward farming is not limited to crypto-related airdrops, and similar ones can be found in “traditional” loyalty programs.
@@
 Given that similar problems exist even in traditional contexts, where users can be easily identified and penalized, it seems that reward farming cannot be solved easily.
-Distributing governance tokens may be dangerous.
-Some platforms airdrop governance tokens, as a means to decentralize their governance process.
-Such tokens allow holders to vote in the platforms’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value (Dotan et al., 2023).
+5.2 Decentralized Governance Threats
+Some protocols airdrop governance tokens, as a means to decentralize their governance process.
+However, distributing governance tokens may be dangerous.
+Such tokens allow holders to vote in the protocols’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value [17].
 There is some empirical evidence that suggests that airdrops of governance tokens may be more successful than airdrops of non-governance tokens.
 In a recent analysis of the market capitalization of crypto tokens performed by Makridis et al.
-(Makridis et al., 2023), the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped.
-But, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
-Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system (Messias et al., 2023a; Feichtinger et al., 2023; Sharma et al., 2023).
-Also, to users that do not have the platform’s best interests in mind, and may even act to the determent of the platform.
-6. Design Guidelines
-The pitfalls outlined earlier, while indeed sobering, may inform platforms embarking on the airdrop journey, and can shed light on the path to potential success.
-Alternative incentives may better sustain user engagement.
-An airdrop’s long-term benefit to a platform can be tricky to measure, given that the potential benefits may be indirect and hard to quantify in advance, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from platforms to users and is thereafter no longer available for other uses.
+[53], the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped.
+However, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
+Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system [56, 22, 68].
+Also, to users that do not have the protocol’s best interests in mind, and may even act to the determent of the protocol.
+5.3 Insider Trading
+It is a problematic occurrence within organizations where individuals exploit privileged information to gain a financial advantage over other users of the protocol.
+This practice is widely recognized as a violation of securities laws in traditional financial markets [6], and engaging in such activities can lead to negative responses from the blockchain community [69, 61, 23].
+When someone within a protocol utilizes privileged information to enhance their profits during an airdrop, it can trigger a backlash from the blockchain community.
+Insider traders have advanced knowledge of the metrics used to determine eligibility and the amount of an airdrop for each address.
+This knowledge empowers them to manipulate the system and make a profit.
+For example, there were claims that a head of growth at AltLayer might have used insider information to make a $200000$ USD profit from their airdrop, although later dismissed as a mere coincidence [69].
+However, such incidents can undermine the trust users place in these protocols.
+This issue also raises concerns about fairness, as some users have access to superior and more accurate information than others.
+Identifying these traders is a challenging problem, making it crucial for protocols to provide extensive information to their users.
+In addition, incentives for blockchain data analytics companies and research groups to conduct post-airdrop data audits can help identify insider traders by analyzing address details and transfer patterns [23].
+For this to happen, data availability is key.
+Therefore, protocols should ensure transparency and encourage thorough analysis to maintain integrity and fairness within the blockchain community.
+6 Design Guidelines
+The airdrop design challenges outlined earlier, while indeed sobering, may inform blockchain protocols embarking on the airdrop journey, and can shed light on the path to potential success.
+6.1 Alternative Incentives may Better Sustain User Engagement
+An airdrop’s long-term benefit to a protocol can be tricky to measure, given that the potential benefits may be indirect and hard to quantify, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from protocols to users and is thereafter no longer available for other uses.
 Furthermore, some costs and effects cannot be easily foreseen, like those arising from distributing governance tokens to users.
@@
 When considering L2s, these discounts can be given on transaction fees.
-Thus, if a user receives this reward, it necessarily has to use the platform again in the future to benefit from the incentive, thus effectively fostering continued user engagement.
-Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the platform, and the platform’s costs go solely towards users actively using the system.
-Nevertheless, while fee discounts are more focused towards incentivizing continued interaction with the platform, they are not without their own challenges.
+Thus, if a user receives this reward, it necessarily has to use the protocol again in the future to benefit from the incentive, thus effectively fostering continued user engagement.
+Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the protocol, and the protocol’s costs go solely towards users actively using the system.
+Nevertheless, while fee discounts are more focused towards incentivizing continued interaction within the protocol, they are not without their own challenges.
 For example, one should carefully design a proper discount mechanism, covering both the eligibility criteria for users, and the height of awarded discounts.
@@
 Another option is to perform multiple airdrops over a relatively long period of time, instead of a single one.
-Target well-known, reputable entities.
+Although this is still susceptible to some of the pitfalls of standard one-shot airdrops, it may serve to ensure the long-term engagement of the community, and thus avoid the decline in user adoption that some experience immediately after an airdrop.
+A more innovative approach was adopted by Blast, which introduced a point-based reward program [55].
+Under this program, users receive rewards based on the points they accumulate through various activities, such as bridging tokens to the protocol (i.e., transferring users’ funds from another protocol) and participating in a referral program where bringing in more users increases their points.
+Interestingly, Blast received 1.1 billion USD in deposits even before its official launch [43].
+This method provides an evaluative metric for users’ contributions to the protocol, structured around a referral-program model.
+Furthermore, the incorporation of innovative distribution mechanisms demonstrates as a critical role in mitigating the presence of sybils within whitelisted addresses eligible for an airdrop.
+For example, Celestia has proposed a distinctive design that employs GitHub commits as a proxy to assess a user’s contribution to the blockchain ecosystem [27].
+Nevertheless, a potential concern emerges in regard to the possibility of users or farmers engaging in the generation of counterfeit activities on GitHub, aiming to exploit other protocols’ airdrops adopting a strategy analogous to Celestia’s.
+Among other metrics, would be to focus on metrics that are resistant to programmable manipulation, adding a layer of difficulty or cost to automated user accounts.
+6.2 Target Well-known Reputable Entities
 Instead of rewarding pseudonymous users, one can target DAOs that are building relevant apps.
-In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs (Foundation, 2023a).
-Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students (Foundation, 2023c) or members of the technical community (Foundation, 2023b) who wish to take part in research and build tools relevant to the platform.
-Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world (Foundation, 2021).
-Proactive oversight and community engagement.
-While the airdrop is ongoing, the platform’s data should be carefully analyzed to verify that no one is attempting to game the system.
-A good example was put forth by Linea, whose team uncovered a vulnerability (Liena, 2023) that allowed users to cheat their incentive mechanism.
+In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs [24].
+Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students [26] or members of the technical community [25] who wish to take part in research and build tools relevant to the protocol.
+Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world [28].
+6.3 Proactive Oversight and Community Engagement
+While the airdrop is ongoing, the protocol’s data should be carefully analyzed to verify that no one is attempting to game the system.
+A good example was put forth by Linea, whose team uncovered a vulnerability [49] that allowed users to cheat their incentive mechanism.
 The prompt discovery prevented cheaters from obtaining over one-third [^2] A dashboard is available on Dune at https://dune.com/rusty378/guildxyz-pins.
-of non fungible tokens (NFTs)offered as incentives.
-A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities (Team, 2023), whether these are actively used or not.
-For example, a disclosure by an AzukiDAO community member allowed the platform to mitigate a vulnerability relatively quickly (MetaSleuth, 2023).
-A platform’s oversight should extend beyond on-chain data.
-In particular, social media is rife with scammers who advertise “fake” airdrops for their platforms, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds.
-Even platforms that are not planning an airdrop are targeted by such ploys.
+of NFTsoffered as incentives.
+A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities [71], whether these are actively used or not.
+For example, a disclosure by an AzukiDAO community member allowed the protocol to mitigate a vulnerability relatively quickly [58].
+A protocol’s oversight should extend beyond on-chain data.
+In particular, social media is rife with scammers who advertise “fake” airdrops for their protocols, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds.
+Even protocols that are not planning an airdrop are targeted by such ploys.
 Actively engaging with the community regarding technical matters can prove beneficial, as well.
-For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar (cygaar, 2023), who found that it could be implemented more cost-effectively.
-Starkware’s governance team have so far performed admirably in this regard (Chevet, 2023).
-The team organized “town hall meeting”-esque discussions and workshops, and actively invited community members to come, participate and voice their opinions.
-These events provide an effective and transparent communication channel, directly reaching active community members.
-Operating transparently and giving the community visibility into the inner-workings of the platform can assure that even if technical issues arise, they are met with understanding by users.
-Rewards should scale with costs.
+For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar [16], who found that it could be implemented more cost-effectively.
+Operating transparently and giving the community visibility into the inner-workings of the protocol can assure that even if technical issues arise, they are met with understanding by users.
+6.4 Rewards Should Scale with Costs
 The ramifications of Goodhart’s law are readily observable in previous airdrops.
-For example, some airdrops attempt to incentive user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the platform in the past.
-Consequently, users can game the system by performing meaningless wash-trading transactions, meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the platform.
+For example, some airdrops attempt to incentivize user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the protocol in the past.
+Consequently, users can game the system by performing meaningless wash-trading transactions [74], meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the protocol.
 This is further compounded by the fact that action-based metrics used by airdrops do not usually account for the actual costs incurred by users for each action.
 For instance, if the number of transactions is used as a metric, one should consider that if transaction fees are low, an airdrop farmer may cheaply perform the required amount of transactions.
-On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop.
+To address this, adopting a quadratic voting system could deter users from artificially inflating transaction volumes, as their rewards would be proportionate to the square root of their activities [11, 48].
+Consequently, the cost to achieve a similar impact would rise, acting as a barrier against manipulation.
+On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop [60].
 To remedy these issues, rewards should scale with the costs that are incurred by users.
-7. Related Work
-The majority of recent papers on airdrops concern themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
+6.5 Data Availability
+Blockchain data is typically publicly accessible but not easily accessible.
+Ensuring data availability is crucial for monitoring airdrop activities and combating Sybil attacks.
+While previous research on airdrops and Sybil detection has primarily focused on single blockchains, our study expands this scope to multiple blockchains.
+Specifically, we collect data from two major Ethereum rollups, Arbitrum and zkSync Era, and gather and label data from significant airdrops, including a major Sybil attack on zkSync Era (the Gemstone “airdrop”).
+Since Sybil detection is increasingly challenging and requires data from multiple chains—given that protocols often reward participants from different chains [27, 4]—data availability is essential for analyzing past airdrops and testing new metrics for future applications.
+However, the rapid growth of blockchain data can make its collection and processing impractical.
+Protocols should ensure their data is available for researchers and companies, helping improve monitoring tools and facilitate learning from the data.
+To support this, we will make our zkSync Era and Ethereum datasets publicly available, enabling others to utilize them for optimizing Sybil detection and airdrop designs.
+7 Related Work
+The majority of recent papers on airdrops concerned themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
 Studies.
@@
 al.
-(Makridis et al., 2023) investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges.
+[53] investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges.
 They find that airdrops of governance tokens, in particular, are positively associated with growth in market capitalization and volume.
@@
 al.
-(Lommers et al., 2023) offer a broad treatment of the different kinds of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based).
-They offer examples of projects like Blur and the pitfalls of their airdrops, together with some advice on airdrop implementation strategies.
-In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success and we provide practical guidance on how to toggle these.
+[52] offer a broad treatment of the different types of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based).
+They offer examples of projects like Blur and the pitfalls of their airdrops, along with some advice on airdrop implementation strategies.
+In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success, and they provide practical guidance on how to toggle these.
 Fan et.
-al.(Fan et al., 2023) study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristic of community members and the effectiveness of airdrop.
-They find that users receive more rewards tend to take positive actions towards the community.
-They summarize several arbitrage patterns and confirm the current detection is not sufficient in screening out airdrop hunters.
+al.[21] study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristics of community members and the effectiveness of airdrop.
+They find that users who receive more rewards tend to take positive actions towards the community.
+They summarize several arbitrage patterns and confirm that the current detection is not sufficient to detect airdrop hunters.
 They use an unsupervised hierarchical clustering method to capture the behaviors of eligible airdrop addresses after receiving rewards and to divide them into interpretable categories.
-Allen (Allen, 2023) examines airdrop design as co-evolutionary process and makes three contributions.
+Other works focused on addressing the problem of Sybil detection, providing valuable insights into identifying airdrop farmers.
+Commonly adopted practices include graph network analysis [21, 51] and machine learning techniques [36, 79, 73, 13, 59].
+The former involves identifying prevalent transfer patterns, such as sequences, radial structures, or a combination of both.
+The latter utilizes approaches like node embedding and employs universal pre-trained transformer encoders like BERT for effective fraud detection [36] or clustering [73, 59].
+Allen [2] examines airdrop design as co-evolutionary process and makes three contributions.
 First, applying a co-evolutionary lens to airdrops, where participants learn and adapt their strategies (e.g., new criteria, on-chain farming) both within and between web3 ecosystems.
@@
 Allen et.
-al.(Allen et al., 2023) consider why projects might choose this unusual and costly means of token distribution.
+al.[3] consider why projects might choose this unusual and costly means of token distribution.
 It considers a selection of high-profile airdrops as case studies between 2014 and 2022.
 This is the first comprehensive analysis of the rationales and mechanisms of Web3 token airdrops.
-We find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralisation of ownership and control of a project (building community, providing regulatory protection, and enhancing security).
+They find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralization of ownership and control of a project (building community, providing regulatory protection, and enhancing security).
 While many commentators describe airdrops as primarily for marketing, they claim that marketing is a weak rationale for airdropping tokens and the evidence of successful marketing-driven airdrops lacking.
-Technical Aspects of Airdrops.
-Frowis et.
+Technical Aspects of Airdrops  Frowis et.
 al.
-(Fröwis and Böhme, 2019) point out that efficient transfers to a large number of recipients in the case of airdrops present issues on Ethereum.
+[31] point out that efficient transfers to a large number of recipients in the case of airdrops present problems on Ethereum.
 They find that cost savings of a factor of two are in fact possible, but require specific provisions in the smart contract implementing the token system.
-Pull-based approaches, which use on-chain interaction with the recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient.
-Total costs are broadly linear in the number of recipients independent of the technical approach.
+Pull-based approaches, which use on-chain interaction with recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient.
+Total costs are broadly linear in the number of recipients, regardless of the technical approach.
 Wahby et.
 al.
-(Wahby et al., 2020) point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities.
-In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely-used user credentials (ECDSA and RSA).
+[75] point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities.
+In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely used user credentials (ECDSA and RSA).
 Private airdrop for RSA builds upon a new zero-knowledge argument of knowledge of the factorization of a committed secret integer, which may be of independent interest.
 Their fastest implementation takes 40–180 ms to generate and 3.7–10 ms to verify an RSA private airdrop signature.
-8. Conclusions
-In this work we conducted analysis for of the common pitfalls of airdrops and propose concrete ways to improve the situation.
-Specifically, we analyze the data pertaining to four recent airdrops and conclude that the majority of proceeds (as high as 95%, typically only within a couple of transfer steps) have been sold via exchanges, implying that the airdrops fall far short of their intended purpose.
-We hope that the suggestions in this paper provide inspiration for researchers and practitioners alike.
-Appendix A Reproducibility
-To ensure complete reproducibility of our empirical analysis, we have openly shared both our dataset and scripts (Messias et al., 2023b).
-The primary data source employed in Section 3 comprises blockchain data extracted from both Ethereum and zkSync Era archive nodes.
-We meticulously collected all blocks, transactions, and events associated with the contracts of the four airdrops under study.
-For Section 4, we sourced data from Growthepie.xyz (Growthepie, 2023).
-This platform provided various metrics (daily transaction count, daily active addresses, median transaction cost, total value locked, fees paid by users, stablecoin market cap) for several key platforms (zkSync Era, OP Mainnet, Arbitrum, Immutable X, Polygon zkEVM).
-All these metrics were smoothed using a 7 day moving average.
-3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">
+8 Conclusions
+In this study, we investigated the prevalent issues with airdrops and developed guidelines to improve their impact.
+Our study analyzed six notable airdrops, uncovering a pattern in which recipients quickly traded their airdropped tokens on exchanges.
+Specifically, 36.51% of ENS tokens, 35.35% of dYdX tokens, and 53.47% of 1inch tokens were exchanged soon after distribution.
+Generally, these tokens were sold after only 1 or 2 transactions, with a median of 2 transactions, demonstrating that airdrops do not promote sustained engagement or attract contributors of significant value.
+Furthermore, our examination of Arbitrum indicated a significant increase in total daily fees during the airdrop event, but a decrease in transaction count per address afterward.
+In particular, other protocols outperformed Arbitrum in the absence of an airdrop.
+The average transaction fees across all protocols followed a similar trend since June 2023, suggesting that airdrops may not be the primary driver of user base growth and can impact activity on other chains.
+We also acknowledge the common airdrop challenges related to airdrop farmers, governance token distribution, and insider trading.
+These challenges serve as valuable lessons for blockchain protocols considering airdrop tokens and offer insight into the potential path to success.
+As a measure to avoid these pitfalls, we propose design guidelines for airdrop campaigns, which are summarized below.
+Alternative incentives may better sustain user engagement.
+An interesting approach involves offering alternative incentives, such as fee discounts for future interactions within the protocol, to foster long-term engagement.
+Additionally, implementing multi-round airdrops allows for learning from past rounds, enabling iterative improvements to address previously identified issues.
+Another approach involves exploring innovative metrics to assess the eligibility of an address for receiving an airdrop.
+This might also be a promising tactic to address concerns related to airdrop farming who typically tries to game the system by utilizing past airdrop actions.
+However, this approach may not completely solve the challenge of insider trading.
+Target well-known reputable entities.
+Prioritizing established and reputable entities, including those building on the protocol, research groups, tech communities, and students, can contribute to sustained engagement.
+Funding these entities or directing airdrops towards DAOs developing relevant dApps presents an opportunity to attract value-driven users and promote long-term involvement.
+Proactive oversight and community engagement.
+Critical to the success of any airdrop is proactive oversight and community engagement.
+Analyzing data from past airdrops, even those on different chains, provides insights into pitfalls and allows for adjustments in future airdrop strategies.
+Open communication channels with the community help understand their interests and desired protocol functionalities.
+Rewards should scale with costs.
+Ensure that rewards scale proportionally with associated costs to prevent exploitation by airdrop farmers.
+Aligning rewards with costs is a strategic measure to mitigate potential gaming of the system.
+In conclusion, our paper offers insights into airdrop challenges and proposes practical guidelines for designing effective airdrop campaigns.
+We hope these suggestions inspire blockchain protocols, researchers, and practitioners.
+Finally, in our commitment to promoting scientific reproducibility, we plan to share the dataset and scripts utilized in this study [57].

### v3

@@
-
+Johnnatan Messias 11    Aviv Yaish 22    Benjamin Livshits 33
 Abstract
-Airdrops are used by blockchain applications and protocols to attract an initial user base, and to grow the user base over time.
-In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying protocol, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop.
-Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
-In this work, we outline the design space for airdrops, and specify a reasonable list of outcomes that an airdrop should ideally result in.
-We then analyze on-chain data from several larger-scale airdrops to empirically evaluate the success of previous airdrops, with respect to our desiderata.
-In our analysis, we demonstrate that airdrop farmers frequently dispose of the lion’s share of airdrops proceeds via exchanges.
-Our analysis is followed by an overview of common pitfalls that common airdrop designs lend themselves to, which are then used to suggest concrete guidelines for better airdrops.
-keywords:
-Airdrop design, Airdrop farmer, Web3, Layer 2
+Airdrops are a common strategy used by blockchain protocols to attract and grow an initial user base.
+Tokens are typically distributed to select users as a “reward” for engaging with the protocol, aiming to foster long-term community loyalty and sustained economic activity.
+Despite their prevalence, there is limited understanding of what makes an airdrop successful.
+This paper outlines the design space for airdrops and proposes key outcomes for an effective strategy.
+We analyze on-chain data from six large-scale airdrops to assess their success and find that a substantial portion of tokens is often sold off by “airdrop farmers.” Based on this analysis, we highlight common pitfalls and offer guidelines for improving airdrop design.
 1 Introduction
@@
 For instance, certain airdrop mechanisms focus on power users, granting them a substantial amount of rewards with the expectation that these users will stimulate valuable economic activity, subsequently attracting additional users.
-However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance.
-It allows users to suggest or vote on changes within a platform, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token.
-This approach introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making voting power [56, 22, 68].
+However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token.
+This introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making power [17, 48, 60].
 To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops.
-Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process.
+We go over previous airdrops, see how they measure up, and uncover some interesting insights in the process, comparing them to a basic desiderata.
 Specifically, we analyze the data pertaining to five popular airdrops (ENS [14], dYdX [33], 1inch [7], Arbitrum [35], and Uniswap [1]) and a fake airdrop conducted by a Sybil farmer (Gemstone [30, 42]).
-We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of airdropped tokens that they receive.
+We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of tokens that they receive.
 Furthermore, we describe the common challenges faced by airdrops in the past.
-Given that the blockchain ecosystem is still quite young, it is expected that the theory and practice of blockchain airdrops are still in its infancy and thus that previous airdrops may not have been entirely successful in the long run.
-Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms that are more fair to honest users.
-1.1 Research Questions
-In this section, we address three critical research questions central to our study.
-[…]
-We evaluate the effectiveness of blockchain protocols in converting short-term user incentives, such as airdrops, into long-term user engagement.
-The primary goal of conducting an airdrop is to attract users in the short-term and encourage them to become long-term users.
-Our investigation aims to measure how well these protocols achieve this objective.
-Contrary to the assumption that airdrops are the primary driver of long-term user base growth, our research reveals otherwise.
-Through an exploration of past airdrops conducted, we demonstrate that airdrops may not be the sole determinant of sustained user retention.
-For example, findings from a recent report by Messari [8] indicate that users join a platform when short-term incentives are available, such as airdrops.
-However, these protocols struggle to retain users in the long-term, as users often shift their attention to protocols offering better incentives as soon as the initial rewards expire.
-This behavior implies that protocols successfully onboard users during incentive periods, like airdrops, but witness a drop in user engagement once the incentives come to an end.
-Additionally, protocols not conducting airdrops also manage to onboard users due to an expectation of potential future airdrops.
-Our findings align with this observation, as our analysis indicates that, even after Arbitrum conducted an airdrop, other protocols attracted user attention and increased their total value locked (TVL).
-This suggests that airdrops alone may not be sufficient to ensure sustained user retention of the platform performing the airdrop, and other factors play crucial roles in long-term platform success.
-[…]
-Since protocols face challenges in achieving sustained user retention through airdrops alone, an alternative strategy involves incentivizing developers and innovators to build new applications, such as decentralized exchanges (DEXs)and non-fungible token (NFT).
-This approach aims to foster a dynamic ecosystem that benefits all users on the platform.
-To enhance user experience and maintain user loyalty, a focus on developing more, potentially exclusive, decentralized applications (dApps) is crucial.
-Another interesting approach is to provide additional incentives, such as discounts on fees for future platform interactions.
-For example, in the context of a rollup, users who receive airdropped tokens could redeem them for discounted fees, encouraging long-term engagement.
-Also, implementing multi-round airdrops allows protocols to learn from previous rounds, enabling iterative improvements to address previously identified issues [2].
-Community engagement also plays a vital role in understanding the needs and preferences of users.
-In line with these considerations, Blast has introduced an innovative airdrop strategy centered around user contributions to their protocol [55].
-This strategy involves awarding points to users as a quantitative measure of their contribution.
-Points are granted for actions such as successfully bringing in new participants to the blockchain protocol, similar to a referral system, or for users who bridge (send) their tokens from one chain to Blast.
-Furthermore, user engagement in blockchain protocols often hinges on short-term incentives.
-Users tend to join the platform when these incentives, such as airdrops, are available.
-However, a notable trend is the propensity of users to migrate opportunistically from one blockchain to another, following the trail of incentives.
-Consequently, a platform experiences an influx of users during incentive periods, only to witness a decline when these incentives cease.
-[…]
-A crucial aspect to consider involves airdrop farmers.
-These are sophisticated users who employ intricate tactics to maximize their chances of receiving more tokens per airdrop.
-They often create numerous wallet addresses controlled by a single entity, a practice known as Sybil [37, 64].
-Detecting Sybils can be challenging [21, 51, 36, 73, 59], especially when these accounts execute automatic and randomized actions (e.g., token transfers, bridging, providing liquidity, exchanging tokens, NFTs, etc.) on the blockchain without linking their wallets.
-Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such actions [70, 29, 12, 15].
-Given its intricacy, current methods for identifying Sybils predominantly rely on graph network analysis [21, 51] and machine learning techniques [36, 73, 59].
-Another common challenge involves airdropping governance tokens.
-These tokens empower holders to propose or vote on protocol changes.
-Distributing these tokens can be risky when users do not prioritize the protocol’s best interests.
-Additionally, it poses the danger of granting excessive power to certain users, as these protocols typically use a one-token-one-vote mechanism [56, 22, 68].
-While acknowledging the challenges highlighted earlier, we believe that these challenges can serve as valuable lessons for protocols venturing into the realm of airdrops, offering insights into the potential path to success.
-Furthermore, it is essential to uphold a positive reputation within the community by avoiding any occurrences of insider trading [5, 42].
-Insider trading refers to individuals within an organization using privileged information to gain a financial edge over other protocol users.
-This practice is a common violation of securities laws in traditional financial markets [6], and engaging in it can lead to negative reactions from the blockchain community [69, 61, 23].
-Finally, the points discussed above raise questions about whether protocols are effectively achieving their goals post their airdrop campaigns.
-1.2 Summary of our Contributions
+Given that airdrops are a relatively recent phenomenon, it is expected that the theoretical and practical understanding of them are still in their infancy, and thus previous airdrops may not have been entirely successful in the long run.
+Finally, we use our analysis to suggest ways to create better airdrop mechanisms that are more fair to honest users.
 We summarize our contributions as follows.
@@
 2 Airdrop Desiderata
-Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with the blockchain protocol or application.
+Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with these protocols and their applications.
 They are widely utilized for these purposes, with numerous examples documented in the literature (see Table 1).
-Any protocol capable of creating tokens can utilize airdrops to distribute them to users.
-For example, blockchain rollup solutions like Arbitrum and Optimism, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
+Protocols can create tokens and use airdrops to distribute them to users.
+For example, blockchain rollup solutions like Arbitrum, Optimism, and ZKsync Era, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
 Airdrops can manifest in various forms, with single-round and multi-round distributions being the most prevalent [2].
@@
 The choice between single-round and multi-round airdrops depends on the protocol’s objectives and community dynamics.
-Community engagement is a critical aspect of airdrop development; it should be carefully considered by protocol developers.
 Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens.
 Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner.
-This difference in user base size introduces an added layer of complexity when performing sybil detection, as there are more accounts to evaluate and potentially safeguard against.
+This difference in user base size introduces complexity when performing Sybil detection, as there are more accounts to evaluate and potentially safeguard against.
 If not performed correctly, this could impact community happiness.
 In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community.
-However, addressing industrial farming poses significant challenges and complexities.
-To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils.
+To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils [39].
 Under this system, sybils have the option to self-report and receive 15% of their designated token allocation.
-Failure to self-report may result in them not receiving tokens if they are identified by the company or its affiliates [46].
 Table 1: List of popular airdrops we analyze in this paper.
+Airdrop Start date End date Blockchain Occurrence Project type
+ENS [14] Oct. 2021 May 2022 Ethereum Single Name service
+dYdX [33] Sept. 2021 Still active Ethereum Continuous DEX
+1inch [7] Dec. 2020 Still active Ethereum Continuous DEX
 […]
-[…]
-dYdX Sept. 2021 Still active Ethereum Continuous DEX
-1inch Dec. 2020 Still active Ethereum Continuous DEX
-[…]
-Arbitrum Mar. 2023 Sept. 2023 Arbitrum One Single Single
-Uniswap Sept. 2020 Still active Ethereum Continuous DEX
-[…]
-Next, we break down airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals.
-These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
+Arbitrum [35] Mar. 2023 Sept. 2023 Arbitrum One Single Single
+Uniswap [1] Sept. 2020 Still active Ethereum Continuous DEX
+Next, we decompose airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals.
+These sub-goals are not necessarily orthogonal and there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
 Attract users in the short-term.
@@
 This can be achieved by issuing rewards that can only be used within the blockchain protocol or application, similarly to frequent-flyer points.
-For example, in layer-2 blockchains (L2s) [67], one may award discounts on future transactions.
-Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the blockchain has to offer.
-For example, Linea’s airdrop quests gave users an in-depth tour of their various features and use-cases [44].
+For example, in Layer-2 blockchains (L2s) [59], one may award fees discounts on future transactions.
+Other measures that may help are having multiple airdrops and rewarding users who perform “quests” that expose them to what the protocol has to offer.
+For example, Linea’s airdrop quests gave users an in-depth tour of their features and use-cases [37].
 Target value-creating users.
 An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform.
-In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [78] and DEXs [77], or across a variety of tokens.
-In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable.
+In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [72] and Decentralized Exchanges (DEXs) [70], or across a variety of tokens.
+In rollups, one may consider so-called “creators” who deploy popular and useful contracts or users who bridge tokens to the rollup [74] as particularly valuable.
 Such users provide additional use cases for a platform, and thus serve to attract other users.
@@
 Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them.
-For example, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them [38], and 61% have sold their tokens shortly after the drop in the case of ParaSwap [21].
-In both instances, the majority of recipients ceased using the relevant blockchain protocols within a few months.
+For example, shortly after the ParaSwap’s drop, 61% of their tokens where sold [16].
+In both cases, the majority of recipients ceased using the relevant blockchain protocols within a few months.
 This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely Sybil accounts.
-Additionally, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects.
-Here, we examine data associated with six airdrops (ENS, dYdX, 1inch, Arbitrum, and Uniswap, including Gemstone) presented in Table 1, which were gathered from archive nodes of Ethereum and zkSync Era.To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query: select * from addresses_ethereum.dex and from Etherscan [20].
-Table 2: General airdrop distribution statistics.
-[…]
-supply distributed exchanges addresses
-[…]
-dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$
-1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$
-Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$
-[…]
-Table 3: Tokens received per address.
-[…]
-[…]
-dYdX $2325.85$ $91336.75$ $392.56$ 00 $24.7\times 10^{6}$
-1inch $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
-Gemstone $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
-Arbitrum $1874.02$ $1566.92$ $1250$ $625$ $10250$
-Uniswap $620.13$ $9347.91$ $400$ $400$ $2.1\times 10^{6}$
-[…]
+Moreover, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects.
+Here, we examine data associated with six airdrops presented in Table 1, which were gathered from archive nodes of Ethereum, Arbitrum, and ZKsync Era (see §0.A for details on our data collection).
+To identify exchanges, we use a list of 620 exchange account addresses obtained from Dune and Etherscan [15].
+Table 2: Airdrop distribution statistics for the six airdrops.
+Note that protocols often send a high share of airdropped tokens to addresses they control.
+For more details on the top earners, refer to Table 4 in §0.D.
+Protocol Token Tokens % sold at # of recipient Tokens received per address
+supply distributed exchanges addresses Mean Median Min Max
+ENS $100\times 10^{6}$ $19.6\times 10^{6}$ 36.62% $102822$ $190.93$ $152.4$ $0.0001$ $1143.54$
+dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$ $2325.85$ $392.56$ 00 $24.7\times 10^{6}$
+1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$ $1917.86$ $636.48$ 00 $9.75\times 10^{6}$
+Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$ $2.31\times 10^{12}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
+Arbitrum $10\times 10^{9}$ $1.09\times 10^{9}$ 10.15% $583137$ $1874.02$ $1250$ $625$ $10250$
+Uniswap $1\times 10^{9}$ $136.5\times 10^{6}$ 5.50% $220220$ $620.13$ $400$ $400$ $2.1\times 10^{6}$
 3.1 Tokens Distributed Over Time
-Table 2 presents the number of airdrops distributed and subsequently sold on exchanges.
-Additionally, Table 3 details the distribution of airdrops per user.
-These tables provide quantitative insights into how airdropped tokens are disseminated across the user base.
+We provide quantitative insights into how airdropped tokens are disseminated across the user base in Table 2.
 Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2).
@@
 It is important to highlight that Gemstone was executed primarily as a Sybil attack rather than a legitimate airdrop.
-Figure 1: Tokens claimed per day for each of the Airdrops analyzed.
+(a) Daily token claims.
+(b) Time to reach an exchange.
+Figure 1: Comparison of token claiming and transfer patterns post-airdrop: (a) daily token claims; (b) time taken to reach an exchange.
 Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users.
-Figure 1 illustrates the distribution of accounts claiming their tokens each day.
+Fig.
+1(a) illustrates the distribution of accounts claiming their tokens each day.
 Arbitrum conducted a substantial airdrop, distributing $1162166000$ ARB tokens among $625143$ selected accounts.
@@
 Table 3: Airdrop recipients’ interaction with exchanges.
-[…]
-Total # sent to exchanges % sent to exchanges
-[…]
-dYdX $42682$ $41320$ 96.81%
-1inch $43402$ $41913$ 96.57%
-Gemstone $20014$ $20000$ 99.93%
-Arbitrum $583137$ $555040$ 95.18%
-[…]
-Table 5: Distribution of hops for airdrop recipients connecting with exchanges.
-[…]
-Mean Std. Median Min. Max.
-[…]
-dYdX 2.25 0.59 2 1 18
-1inch 1.85 0.64 2 1 7
-Gemstone 1 1 1 1 1
-Arbitrum 2.71 $0.79$ $3$ $1$ $32$
-Uniswap 2.11 $0.43$ 2 1 9
-[…]
-(a) Over the number of post-airdrop blocks.
-(b) Over the number of post-airdrop days.
-Figure 2: Time for airdrop tokens to be transferred to exchanges.
+Gemstone and 1inch stands out as an exception with a median of 1 hop, while other protocols show a median of 2 hops for connecting airdrop recipients with exchanges.
+Protocol Airdrop recipients # of hops to reach an exchanges
+Total # sent to exchanges % sent to exchanges Mean Median Min. Max.
+ENS $102822$ $86156$ 83.79% 1.65 2 1 50
+dYdX $42682$ $40475$ 94.83% 1.77 2 1 17
+1inch $43402$ $41913$ 96.57% 1.50 1 1 7
+Gemstone $20014$ $20000$ 99.93% 1 1 1 1
+Arbitrum $583137$ $555036$ 95.18% $2.34$ $2$ $1$ $31$
+Uniswap $220220$ $213066$ 96.75% 2.21 2 1 17
 Users typically interact with exchanges to convert one token for another and to sell them.
@@
 This engagement ranges from a minimum of 83.79% in the case of ENS to a maximum of 99.93% for Gemstone.
-Additionally, Table 5 presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset.
+Additionally, Table 3 also presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset.
 Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities.
 Gemstone is a notable outlier, with all tokens being sold within a single hop.
-Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (between 2 and 3 hops).
+Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (2 hops).
 This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
+The majority of accounts exchanged their tokens within approximately one million blocks post-airdrop.
+The number of blocks for Gemstone is significantly higher than others due to a delay introduced by its developer.
+Given the difference in block generation times between Ethereum (with a new block created every 15 seconds) and ZKsync, we normalized the blocks in terms of days.
+From Fig.
+1(b), within a single day, 66.09% of the 1inch accounts interacted with an exchange.
+In comparison, ENS had an interaction rate of 55.15%, while dYdX stood at 64.26% followed by Arbitrum and Uniswap with 60.34% and 12.39%, respectively.
+This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in §2.
+The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
+3.2 Token Distribution Graphs
+To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network.
+Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another.
+Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20014$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges.
+Lastly, Arbitrum accounts for $2025898$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
+To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Fig.
+7 in §0.B.
+We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [15].
+We show that for all protocols but Gemstone, the DEXthat received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.
+In the case of Gemstone, all tokens were sent to the address 0x7$\cdots$49ad in just a single hop.
+On the other hand, in dYdX, the airdrop utilized a more diverse set of exchange addresses.
+Notably, as depicted in Table 3, a portion of the airdrop recipients chose to sell their tokens on an exchange.
+We see some common exchanges like Uniswap, Wintermute, and SushiSwap.
+(a) Absolute number.
+(b) Relative daily active addresses.
+Figure 2: Unique daily active addresses per protocol: (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
+(a) Daily median fees.
+(b) Daily average fees.
+Figure 3: Daily transaction fees in USD. (a) Median; and (b) Average.
+4 Measuring the Airdrop Lift
+Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [16].
+Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [16, 25].
+In this section, we examine this through the lens of some relevant metrics such as daily transaction count, daily active addresses, median transaction cost, Total Value Locked (TVL), fees paid by users, and stablecoin market cap applied to the Arbitrum airdrop.
+We rely on data source from Growthepie [29].
+For more details on this data refer to §0.A in the appendix.
+4.1 Unique Active Addresses
+Others outperformed Arbitrum without an airdrop.
+While the number of unique addresses on Arbitrum increased post-airdrop (March 22nd 2023 [57]) and remained 50% higher than pre-airdrop levels, other protocols achieved similar growth without an airdrop.
+For example, per Fig.
+2(b), Optimism saw a larger increase in addresses in May 2023, likely due to the Bedrock launch [54].
+Similarly, ZKsync Era surpassed Arbitrum’s address count two months after the airdrop, as shown in Fig.
+2(a).
+Fees may explain the closing gap between Arbitrum and Optimism.
+Arbitrum has consistently led Optimism in daily active addresses.
+However, Fig.
+8 in §0.C shows this gap narrowing.
+Before the airdrop, Arbitrum had 2.6 times more active addresses than Optimism, but in the last 50 days, this has decreased to 1.83 times.
+Optimism’s lower median transaction fees since June may partly explain this trend (Fig.
+3(a)), with the average fee gap between the two protocols being even smaller (Fig.
+3(b)).
+Unique address counts oscillates over time.
+The number of unique addresses shows oscillatory behavior, as seen in Fig.
+2(b), with rapid rises, peaks, and drops.
+Notably, Optimism and Arbitrum display inverted phases in relative address counts, possibly due to users switching between protocols as fees rise.
+However, Fig.
+3(a) shows no such pattern in median transaction fees, with Arbitrum consistently having lower fees in mid-May 2023.
+However, unique address counts metric can be gameable.
+Per our analysis, Arbitrum’s airdrop did not lead to long-lasting user involvement, as the unique address metric can be gamed.
+Users can create multiple addresses (Sybil attacks) to exploit airdrop limits.
+This makes the metric unreliable for measuring genuine activity, as large fluctuations may result from such behavior.
+Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such activities [10, 24, 53, 63].
+Thus, alternative, more Sybil-resistant metrics should be considered to gauge authentic user involvement such as graph network analysis [16, 43] combined with machine learning techniques [31, 50, 66].
+4.2 Transaction-related Metrics
+Transaction-related metrics provide a useful proxy for measuring “real” economic activity, as users must pay a fee to send transactions [27], excluding cases with retroactive rebates or protocols operating at a loss [55].
+The Arbitrum-Optimism gap narrows when considering transactions.
+Notably, the transaction count gap between Arbitrum and Optimism had nearly closed by the end of July (see Fig.
+4(a)).
+Additionally, Fig.
+4(b) shows that Immutable X’s daily transaction count has almost halved since Arbitrum’s airdrop, whereas its unique address count remained relatively stable.
+This suggests that while address numbers held steady, user engagement with Immutable X has decreased.
+(a) Absolute daily transactions.
+(b) Relative daily transactions.
+Figure 4: Daily number of transactions per protocol. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
+Arbitrum’s per-address transaction count declined post-airdrop.
+To assess how Arbitrum’s airdrop impacted user engagement in other protocols, Fig.
+5(a) shows the relative average daily transaction count per unique address.
+Since the airdrop, Arbitrum’s per-user transaction count has dropped to less than 75% of its pre-airdrop level.
+However, since transaction counts can be misleading without considering fees, high transaction counts alone may not indicate genuine engagement.
+In this regard, some protocols require users to perform multiple transactions to qualify for airdrops, leading to inflated activity when fees are low.
+This aligns with Goodhart’s law [46], which states that “when a measure becomes a target, it ceases to be a good measure”.
+(a) Daily transactions per address.
+(b) Average fees spent per address.
+Figure 5: Daily transactions per address (a) and daily average fees per address (b), relative to the average before Arbitrum’s airdrop.
+Since June, average transaction fees are similar for all protocols.
+Good metrics should reflect a user’s commitment, and transaction fees can serve as a proxy for this since it measures how much a user is willing to expend for interacting with the protocol.
+Since June, the average fee per transaction and per unique address have been similar across protocols.
+Comparing average and median fees (Fig.
+3) shows that the median fee may be more useful in understanding shifts in user behavior.
+Another useful metric is the relative average fees per address compared to the 50 days before Arbitrum’s airdrop, as shown in Fig.
+5(b).
+It shows that Arbitrum’s user engagement was not significantly impacted by the airdrop and generally follows similar patterns to other protocols.
+Arbitrum’s total daily fees spiked around the time of the airdrop.
+Arbitrum’s airdrop did not provide a significant long-term advantage over its competitors in terms of income from transaction fees as shown in Fig.
+9 in §0.C.
+While Arbitrum experienced a peak in fees on the day of the airdrop, this increase was short-lived.
+Indeed, per Fig.
+6(a), in the 50 days before the airdrop, Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
+(a) Ratio of daily transaction fees.
+(b) TVL growth.
+Figure 6: Comparison of the ratio of daily transaction fees between Arbitrum and Optimism (a); and the growth in Total Value Locked (TVL) relative to the average of the 50 days before Arbitrum’s airdrop (b).
+4.3 Total Value Locked (TVL)
+The TVLof a protocol is a metric that measures the combined value of all assets stored within that protocol.
+[…]
+5 Common Airdrop Design Challenges
+Airdrops, much like traditional loyalty programs—such as newcomer bonuses provided by banks and credit card companies—face several common design challenges.
+However, the unique context of blockchain technology and the specific mechanics employed by most airdrops can amplify these challenges or even introduce new ones.
+In this section, we explore three of these challenges.
+5.1 Airdrop Farmers
+These are sophisticated users who employ elaborate tactics to maximize the amounts of airdrop tokens they receive.
+Blockchain protocols have adopted various measures to mitigate airdrop farmers gaming the system.
+A popular approach is to limit the amount of rewards that a single user can receive.
+Thus, protocols resort to Proof-of-Humanity (PoH)services, such as Gitcoin passport [28].
+These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human.
+Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH [28].
+Such methods can be augmented by analyzing on-chain data to detect Sybils and exclude them from the airdrop [5], but this may lead to false negatives [69].
+There are some recent examples of protocols like Celestia using Sybil resistance solutions [40].
+There are other mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions [56], to sharing posts on social media [38].
+These tasks sometimes appear arbitrary, leading to frustration among users and they are easy to automate [34] and cheaply fooled, even more-so when protocols offer retroactive rebates on transaction fees [55].
+Furthermore, the reliance of many protocols on a limited number of PoHservices means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops, even with biometric identification, full Sybil resistance is not guaranteed [65].
+A different approach used by protocols is to announce airdrops that retroactively reward users that were active before the announcement.
+Nevertheless, farmers can prepare in advance by interacting with these protocols even when no airdrop has officially been announced [9], as was the case in dYdX’s airdrop [13].
+The phenomena of reward farming is not limited to crypto-related airdrops, and similar ones can be found in “traditional” loyalty programs.
+In particular, the practice of credit card churning, where users apply for credit cards for the sole purpose of receiving the rewards bestowed upon newcomers, only to cancel the cards once the rewards were received.
+Given that similar problems exist even in traditional contexts, where users can be easily identified and penalized, it seems that reward farming cannot be solved easily.
+5.2 Decentralized Governance Threats
+Some protocols distribute governance tokens through airdrops as a means to decentralize their governance processes.
+However, distributing governance tokens can be risky.
+These tokens allow holders to participate in the protocol’s governance by voting on key decisions.
+Typically, these tokens can also be exchanged for other tokens, potentially giving them monetary value [12].
+This can lead to more farmers farming these tokens.
+Empirical evidence suggests that airdropping governance tokens may be more effective than those of non-governance tokens.
+In a recent analysis, Makridis et al.
+[45] found that airdropped governance tokens experienced a growth rate in market capitalization up to $14.99$% higher compared to governance tokens that were not airdropped.
+However, the authors also note that this effect is statistically insignificant when using common benchmarks.
+Despite these potential benefits, airdropping governance tokens poses significant risks if not done well.
+It can concentrate too much power in the hands of a few users, leading to an inequitable distribution of decision-making power within the system [17, 48, 60].
+Additionally, some recipients may not have the protocol’s best interests in mind and could act in ways detrimental to the protocol’s long-term success by voting on changes to the protocol that benefit themselves.
+5.3 Insider Trading
+This issue arises when individuals exploit privileged information for financial gain at the expense of other protocol users.
+Such practices are widely recognized as violations of securities laws in traditional financial markets [4] and often provoke negative reactions from the blockchain community [18, 52, 61].
+When someone within a protocol uses privileged information to increase their profits, it can trigger a community backlash.
+Insiders may have advanced knowledge of the metrics used to determine eligibility and the reward for each address, and may use it to their advantage.
+For example, there were claims that a head of growth at AltLayer might have used insider information to make a $200$K USD profit from their airdrop, although later dismissed as a mere coincidence [61].
+However, such incidents can undermine the trust users place in these protocols.
+This issue also raises concerns about fairness, as some users have access to superior and more accurate information than others.
+Identifying these inside traders is a challenging problem, making it crucial for protocols to provide extensive information to their users.
+In addition, incentives for blockchain data analytics companies and research groups to conduct post-airdrop data audits can help identify insider traders by analyzing address details and transfer patterns [18].
+For this to happen, data availability is key.
+Therefore, protocols should ensure transparency and encourage thorough analysis to maintain integrity and fairness within the blockchain community.
+6 Design Guidelines
+The design challenges outlined earlier, while sobering, may inform future airdrop designers, and can shed light on the path to potential success.
+6.1 Alternative Incentives for Sustaining User Engagement
+The long-term benefits of an airdrop to a protocol can be difficult to measure, as the potential advantages may be indirect and hard to quantify, while the expenses are often immediate and irrevocable.
+Additionally, some costs and effects, such as those arising from distributing governance tokens, may be unpredictable.
+Therefore, instead of using airdrops, communities could consider alternative measures with a more predictable relationship between costs and benefits.
+One straightforward alternative is for the community to vote to programmatically reward loyal users with discounts on future interactions.
+In the context of Layer 2 (L2) solutions, these discounts could be applied to transaction fees.
+This approach encourages users to engage with the protocol again to benefit from the incentive, thereby fostering continued user participation.
+Additionally, this incentive mechanism is somewhat more resistant to airdrop farming, as the discounts have no intrinsic value outside the protocol, and the protocol’s costs are directed solely toward users who actively use the system.
+However, careful design of the discount mechanism is essential, including determining eligibility criteria for users and setting appropriate discount levels.
+Moreover, it is not immediately clear whether discounts can attract users as effectively as the more immediate and tangible rewards offered by airdrops.
+Another option is to perform multiple airdrops over an extended period instead of a single event.
+Although this approach remains susceptible to some pitfalls of standard one-shot airdrops, it may help ensure long-term community engagement and prevent the decline in user adoption that some protocols experience immediately after an airdrop.
+A more innovative approach was adopted by Blast, which introduced a point-based reward program [47].
+Under this program, users earn rewards based on points accumulated through various activities, such as bridging tokens to the protocol (i.e., transferring funds from another protocol) and participating in a referral program that increases points for bringing in more users.
+Notably, Blast received \$1.1 billion in deposits even before its official launch [36].
+This method provides a measurable metric for users’ contributions to the protocol, structured around a referral-program model.
+Furthermore, the incorporation of innovative distribution mechanisms plays a crucial role in mitigating the presence of sybil attacks within whitelisted addresses eligible for an airdrop.
+For example, Celestia has proposed a unique design that uses GitHub commits as a proxy to assess a user’s contribution to the blockchain ecosystem [22].
+However, a potential concern arises with the possibility of users or farmers generating counterfeit activities on GitHub to exploit other protocols’ airdrops adopting a similar strategy to Celestia’s.
+Farmers, therefore, might expect that new protocols use similar selection criteria than past airdrops.
+To counter this, protocols could focus on metrics that are resistant to programmable manipulation, thereby adding a layer of difficulty or cost to the creation of automated user accounts.
+6.2 Targeting Well-Known Reputable Entities
+Instead of rewarding pseudonymous users, protocols can target developers and projects that are building relevant applications.
+For example, in Arbitrum’s airdrop, 1.13% of the distributed tokens were allocated to DAOs projects [19].
+Arbitrum also offers additional incentives beyond its airdrop for specific groups, such as university students [21] and members of the technical community [20] who wish to participate in research and develop tools relevant to the protocol.
+Another approach was implemented by Optimism, which allocated some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup to the blockchain world [23].
+Prioritizing established and reputable entities, including those building on the protocol, research groups, tech communities, and students, can promote sustained engagement.
+By funding these entities, protocols may attract value-driven users and foster long-term involvement.
+6.3 Proactive Oversight and Community Engagement
+During the airdrop process, continuous monitoring and analysis of the protocol’s data are crucial to prevent exploitation.
+A notable example is provided by Linea, whose team identified a vulnerability that allowed users to manipulate their incentive mechanism [41].
+This prompt discovery averted cheaters from claiming over one-third of the Non-Fungible Tokens (NFTs)offered as incentives [58].
+In addition to technical oversight, engaging the community by maintaining open communication channels and offering bug bounties encourages the disclosure of vulnerabilities, whether exploited or not [64].
+For instance, an AzukiDAO community member disclosed a vulnerability, allowing the protocol to quickly address it [49].
+Oversight should also extend beyond on-chain data.
+For example, social media is frequently exploited by scammers promoting fake airdrops, luring users into connecting their wallets to fraudulent websites that drain their funds.
+Even protocols not planning an airdrop may be targeted by such schemes.
+Active community involvement in technical discussions is also advantageous.
+For example, ZKsync Era’s NFTdrop was retrospectively analyzed by cygaar [11], who identified potential cost-saving improvements.
+Maintaining transparency and providing the community with insights into the protocol’s inner workings fosters trust.
+When technical issues arise, users are more likely to respond with understanding when they are well-informed.
+6.4 Rewards Should Scale with Costs
+The effects of Goodhart’s law [46] are evident in many past airdrops.
+For example, airdrops often aim to incentivize user engagement by publicly announcing that eligibility for rewards is based on the frequency of past interactions with the protocol.
+This approach, however, can be exploited, as users may engage in meaningless wash-trading transactions [67], undermining the criteria as a true reflection of genuine engagement.
+The problem is further exacerbated by the fact that action-based metrics used to determine eligibility often fail to consider the actual costs users incur for each action.
+For instance, when the number of transactions is the primary metric, low transaction fees enable airdrop farmers to cheaply meet the transaction volume required for rewards.
+A potential solution is the adoption of a reputation-based mechanism [32], which could disincentivize artificial inflation of transaction volumes.
+However, the protocol must carefully define what constitutes “user reputation” and determine the appropriate metrics for evaluation.
+Conversely, high transaction fees can erode the value of the rewards users receive, reducing the attractiveness of the airdrop [51].
+To address these issues, rewards should scale with the actual costs incurred by users, ensuring a fairer and more effective distribution of incentives.
+7 Related Work
+Recent research on airdrops has primarily focused on post-factum analyses and guidance for designing effective airdrop campaigns.
+Airdrop Studies  Yaish and Livshits [71] present a theoretical model for airdrops that accounts for both honest users and farmers, where the latter have lower airdrop eligibility costs and lower intrinsic utility from using the platform that issues the airdrop.
+In their analysis, they show that airdrops in which the issuer expends a non-zero fixed cost for each recipient, the threat posed by farmer sybil attacks lead to unbounded issuance costs.
+However, they show losses from farmers can be bounded by setting in advance the total quantity of airdrop tokens and splitting them equally among recipients.
+Moreover, by designing the airdrop mechanism correctly, farmers can be harnessed to foster network effects, thereby attracting honest users who may otherwise would prefer competing platforms.
+Makridis et al.
+[45] explore the impact of governance token airdrops on DEXgrowth.
+Analyzing 51 exchanges, they found that such airdrops significantly contribute to increased market capitalization and trading volume.
+Lommers et al.
+[44] provide a comprehensive overview of various airdrop types, such as plain vanilla, holder-based, and value-based models.
+Their work highlights how eligibility criteria, signaling, and implementation strategies influence the success of airdrop campaigns, offering practical recommendations for optimization.
+Fan et al.
+[16] present a case study of ParaSwap DEX, proposing a role taxonomy methodology based on user behavior and airdrop effectiveness.
+Their findings show that users receiving higher rewards are more likely to contribute positively to the community.
+They also identify arbitrage patterns and note the limitations of current methods for detecting airdrop hunters.
+On the other hand, Sybil detection techniques, such as graph network analysis [16, 43] and machine learning approaches [8, 31, 50, 66, 73], have been proposed to address these issues.
+Allen [2] conducts nine airdrop case studies (e.g., Optimism, Arbitrum, Blur) and offers insights into designs like task-based claiming.
+The study emphasizes the need for more dynamic designs with feedback loops and suggests that some projects may revert to simpler designs due to the complexity and cost of advanced mechanisms.
+Allen et al.
+[3] investigate the motivations behind token airdrops, focusing on marketing and decentralization.
+While airdrops are often seen as marketing tools, the authors argue that this rationale is weak, as evidence of marketing-driven airdrop success is limited.
+Instead, decentralization and community-building are highlighted as the primary motivations.
+Technical Aspects of Airdrops  Frowis et al.
+[26] identify the operational challenges and costs associated with large-scale airdrops on Ethereum.
+They suggest that while cost savings of up to 50% are possible with specific smart contract optimizations, pull-based approaches can shift the cost burden to recipients.
+Overall, the total cost remains proportional to the number of recipients.
+Wahby et al.
+[68] address the privacy issues in current airdrop mechanisms, which leak information about recipients.
+They propose a private airdrop scheme based on zero-knowledge proofs using RSA credentials, achieving privacy while maintaining computational efficiency.
+Their implementation significantly improves both signature generation and verification times.
+8 Conclusion
+This study identified common issues with airdrops and proposed guidelines to improve their effectiveness.
+Our analysis of six major airdrops revealed a consistent pattern of recipients quickly selling their tokens—36.62% of ENS, 35.45% of dYdX, and 54.05% of 1inch tokens were traded shortly after distribution, with a median of just two transactions.
+This suggests airdrops fail to sustain long-term engagement or attract valuable contributors.
+For Arbitrum, we observed a spike in daily fees during the airdrop, followed by a decline in transactions per address.
+Other protocols outperformed Arbitrum without an airdrop, and since June 2023, transaction fees have been similar across protocols, indicating airdrops are not the main driver of user growth.
+Lastly, we discussed challenges like airdrop farming, governance token distribution, and insider trading, offering insights for future airdrop strategies.
+Acknowledgments
+This research article is a work of scholarship and reflects the authors’ own views and opinions.
+It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer.
+Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
+Appendix 0.A Reproducibility
+To ensure full reproducibility of our empirical analysis, we will openly share both our dataset and scripts.
+The primary data source used in §3 consists of blockchain data extracted from Ethereum, Arbitrum, and ZKsync Era archive nodes.
+We collected all relevant blocks, transactions, and events tied to the contracts of the six airdrops examined.
+For §4, data was sourced from Growthepie.xyz [29] corresponding to the year 2023, which offers various metrics, including daily transaction counts, daily active addresses, median transaction costs, TVL, user fees, and stablecoin market capitalization.
+These metrics are available for major blockchains such as ZKsync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM.
+To enhance clarity, all metrics were smoothed using a 7-day moving average.
+Appendix 0.B Graph Network of the Token Distribution
+We construct a graph network (V, E) of the token transfer network for the six airdrops analyzed in this paper and present them in Fig.
+7.
+Each node (V) represents an address, and an edge (E) is created whenever tokens are transferred from one address to another.
+For the ENS network, there are $184585$ nodes and $608462$ edges.
+The dYdX network consists of $112853$ nodes and $406027$ edges.
+The Gemstone network contains $20003$ nodes and $240113$ edges.
+The 1inch network includes $308329$ nodes and $1400913$ edges.
+The Arbitrum network has $921650$ nodes and $27438608$ edges, while the Uniswap network contains $1180830$ nodes and $3762613$ edges.
 (a) ENS.
@@
 Figure 7: Network structure for selling airdropped tokens in 6 airdrops.
-Figure 2 shows the cumulative distribution functions (CDFs) demonstrating how quickly users interact with an exchange right after receiving tokens from an airdrop.
-As shown in Figure 2(a), the majority of accounts exchanged their tokens within approximately one million blocks.
-The number of blocks for Gemstone is significantly higher than others due to a delay introduced by airdrop farmers.
-Given the difference in block generation times between Ethereum (with a new block created approximately every 15 seconds) and zkSync, we normalized the blocks in terms of days.
-According to Figure 2(b), within a single day, 64.67% of the accounts interacted with an exchange.
-In comparison, ENS had an interaction rate of 45.86%, while dYdX stood at 37.41%.
-This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2.
-The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
-3.2 Token Distribution Graphs
-To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network.
-Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another.
-Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges.
-Lastly, Arbitrum accounts for $921650$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
-To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Figure 3.
-We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [20].
-We show that for all protocols but Gemstone, the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.
+Appendix 0.C More Insights Into Airdrop Metrics
+In this section, we provide a complementary analysis of the measures described in Section §4.
+To summarize, Fig.
+8 shows the ratio of unique daily active addresses between Arbitrum and Optimism.
+Additionally, Fig.
+9 presents the total daily transaction fees before and after the Arbitrum airdrop.
+Figure 8: Ratio between the number of unique daily active addresses in Arbitrum and Optimism.
+(a) Absolute daily transaction fees.
+(b) Relative to average.
+Figure 9: Total daily transaction fees before and after the Arbitrum airdrop: (a) absolute values of daily fees in USD; and (b) fees relative to the average before the airdrop.
+Appendix 0.D Top Airdrop Earners
+In this section, we present the top-1 airdrop earners for each of the six protocols analyzed.
+Airdrops can be distributed to various types of accounts, including smart contract accounts, such as the protocol’s own treasury, or Externally Owned Accounts (EOAs).
+Additionally, tokens can be sent to multi-signature (multisig) contract wallets.
+A multisig contract requires a minimum number of users to sign a transaction before it can be executed on-chain [48].
+These multisig accounts may also belong to the protocol itself.
+For instance, the treasury account could receive a substantial portion of the airdropped tokens, which could be used to support the protocol’s ecosystem.
+Therefore, the accounts who receive most of the tokens (the top earners) may be protocol-related accounts (see Table 2 in §3.1).
+To identify the top airdrop earner for each protocol, we list them in Table 4.
+The labels for these earners were obtained using the methodology detailed in §3.2, which involves gathering data from Etherscan, a widely-used blockchain explorer [15].
+We note that analytico.eth alone received $11.71$% of the 1inch airdropped tokens.
+The true identity of the entity controlling this address remains unknown, and it is unclear whether it is managed by a single individual or a group.
+On the other hand, $13.95$% of dYdX were sent to dYdX treasure.
+This is an expected behavior for some airdrops.
+Table 4: Top airdrop earner for the six protocols analyzes in this paper: ENS, dYdX, 1inch, Gemstone, Arbitrum, and Uniswap.
+Protocol Wallet address Label # of tokens % of total supply % of dist. tokens
+ENS 0xa5b5$\cdots$6c2d bentobox.eth $1143.54$ $0.0011$ $0.006$
+dYdX 0x95$\cdots$0b54 TreasuryMerkleClaimProxy $24.7\times 10^{6}$ $2.47$ $13.95$
+1inch 0xa0f7$\cdots$9b73 analytico.eth $9.75\times 10^{6}$ $0.65$ 11.71
+Gemstone 0$\cdots$3a3f Sybil [30] $2.64\times 10^{12}$ $5.21\times 10^{-39}$ $0.005$
 […]
-(a) Absolute number.
-(b) Relative daily active addresses.
-Figure 4: Unique daily active addresses per protocol: (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
-Figure 5: Ratio between the number of unique daily active addresses in Arbitrum and Optimism.
-(a) Daily median fees.
-(b) Daily average fees.
-(c) Average fees spent per unique address per day.
-Figure 6: Daily transaction fees in USD. (a) Median; (b) Average; and (c) Per unique address.
-4 Measuring the Airdrop Lift
-Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [21].
-Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [30, 21].
-In this section, we examine this through the lens of some relevant metrics applied to the Arbitrum airdrop.
-We rely on data source from Growthepie [34].
-This platform supplies various metrics, including daily transaction count, daily active addresses, median transaction cost, TVL, fees paid by users, and stablecoin market cap, for several key protocol such as zkSync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM.
-To enhance clarity, all these metrics were subjected to a 7-day moving average smoothing process.
-4.1 Unique Active Addresses
-Figure 4 examines the number of unique addresses that were active on each protocol using two different approaches.
-The Figure 4(a) depicts the absolute number of addresses.
-The Figure 4(b) shows the relative amount of unique daily active addresses for each protocol, when compared to the same protocol’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 [66].
-This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users [7].
-{observation}
-others outperformed Arbitrum, without an airdrop.
-The absolute number of addresses has increased for Arbitrum post-airdrop, and remained at least 50% higher than its pre-airdrop level in the four months since then.
-However, by looking at other protocols, it becomes apparent that a similar result can be attained without an airdrop.
-This is demonstrated by Optimism’s trajectory in Figure 4(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 [62].
-This is further exemplified by zkSync Era overtaking Arbitrum’s in Figure 4(a), two months after the airdrop.
-{observation}
-fees may explain the closing gap between Arbitrum and Optimism.
-Arbitrum leads over Optimism with respect to the number of unique daily active addresses.
-We turn our focus to Figure 5, which depicts Arbitrum’s lead over Optimism with respect to the number of unique daily active addresses.
-While in the 50 days before the airdrop Arbitrum has enjoyed an average of 2.6 times more unique daily active addresses than Optimism, this gap has recently started to close: the average over the last 50 days in our dataset is considerably lower, and is equal to 1.83 times more.
-Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 6(a).
-We note that with respect to the average fee, the gap between the protocols is even smaller over the same period as shown in Figure 6(b).
-{observation}
-the number of unique addresses oscillates over time.
-Some of the previous observations exhibit oscillatory behavior, in particular in Figure 4(b) depicting the relative amount of unique addresses: rising, reaching a peak, and then subsequently dropping, all in rapid succession.
-In particular, the phases of Optimism and Arbitrum are inverted regarding the relative amount of unique addresses.
-One may attempt to explain this by reasoning that when a peak is reached in one of the protocols, fees should increase due to the load, thus prompting cost-conscious users to turn to a different protocols.
-But, when considering the median transaction fees paid per day (per Figure 6(a)), it is apparent that fees do not have corresponding peaks and dips.
-In fact, Arbitrum consistently had lower median fees in mid-May.
 […]
-4.2 Transaction-related Metrics
-A natural next step is to examine transaction-related metrics.
-In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee [32], when ignoring retroactive rebates or the possibility of blockchain protocols operating at a loss [63].
-{observation}
-[…]
-Some notable differences are that with respect to transaction counts, the gap between Arbitrum and Optimism has nearly closed entirely by the end of July.
-Furthermore, Figure 7(b) shows that the daily transaction count for Immutable X has almost halved since Arbitrum’s airdrop.
-Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the protocol.
-(a) Absolute daily transactions.
-(b) Relative daily transactions.
-Figure 7: Daily number of transactions per protocol. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
-{observation}
-Arbitrum’s per address transaction count dropped since the airdrop.
-To better understand how Arbitrum’s airdrop affected user engagement in other protocols, we now present the Figure 8.
-Figure 8(b), which depicts the relative average daily transaction count per unique address.
-The figure shows that since the drop, Arbitrum’s per-user transaction count has fallen to less than 75% of its pre-airdrop level.
-(a) Average number.
-(b) Relative to average before Arbitrum’s airdrop.
-Figure 8: Daily transactions per address.
-[…]
-{observation}
-since June, average transaction fees are similar for all protocols.
-Conceptually, good metrics should measure a user’s amount of skin in the game.
-A possible proxy for that is the amount of funds the user is willing to expend for interacting with the protocol.
-When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all protocols are quite similar in both regards.
-Furthermore, when comparing the average fee per transaction to the median one (per Figures 6(a) and 6(b)), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
-Another interesting metric is the relative average fees per address, compared to the average in the 50 days preceding Arbitrum’s airdrop, shown in Figure 9.
-One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other protocols.
-Figure 9: Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
-(a) Total daily transaction fees (USD).
-(b) Relative to average before Arbitrum’s airdrop.
-Figure 10: Total fees before and after the Arbitrum airdrop.
-Figure 11: Ratio between the daily fees spent in Arbitrum and Optimism.
-{observation}
-Arbitrum’s total daily fees spiked around the time of the airdrop.
-Figure 10 shows that Arbitrum’s airdrop did not confer a substantial advantage over its competitors when it comes to its income stream from transaction fees.
-Although the total amount of fees it earned on the day of the drop reached a peak, this was momentary.
-Indeed, per Figure 11, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
-(a) Daily TVL, in USD.
-(b) Relative to the average prior Arbitrum’s airdrop.
-Figure 12: Total Value Locked (TVL) growth: (a) Daily TVL in USD; and (b) relative to the average for the 50 days before Arbitrum’s airdrop.
-4.3 Total Value Locked
-The Total Value Locked (TVL) of a protocol is a metric that measure the combined value of all assets stored in that protocol.
-{observation}
-[…]
-5 Common Airdrop Design Challenges
-Airdrops share similarities with traditional loyalty programs, such as newcomer bonuses offered to new customers by banks and credit card companies.
-Furthermore, they also share some common design challenges.
-The blockchain setting and the mechanics adopted by most airdrops may exacerbate these pitfalls, or introduce new ones.
-5.1 Airdrop Farmers
-Sophisticated users known as airdrop farmers employ elaborate tactics to maximize the amounts of airdrop tokens they receive.
-Blockchain protocols have adopted various disproportionate measures to curb the possibility of airdrop farmers gaming the system.
-For example, a popular approach is to limit the amount of rewards that a single user can receive.
-Thus, protocols resort to Proof-of-Humanity services, such as Gitcoin passport [33].
-These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human.
-Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH [33].
-Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop [7], but this may lead to false negatives [76].
-There are some recent examples of protocols like Celestia considering using sybil resistance solutions [47].
-There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions [65], to sharing posts on social media [45].
-These tasks sometimes appear arbitrary, leading to frustration among users.
-The problem with the aforementioned methods is that they are easy to automate [40] and cheaply fooled, even more-so when protocols offer retroactive rebates on transaction fees [63].
-Furthermore, the reliance of many protocols on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops, even with biometric identification, full sybil resistance is not guaranteed [72].
-A different approach used by protocols is to announce airdrops that retroactively reward users that were active before the announcement.
-Nevertheless, farmers can prepare in advance by interacting with these protocols even when no airdrop has officially been announced [14], as was the case in dYdX’s airdrop [18].
-The phenomena of reward farming is not limited to crypto-related airdrops, and similar ones can be found in “traditional” loyalty programs.
-In particular, the practice of credit card churning, where users apply for credit cards for the sole purpose of receiving the rewards bestowed upon newcomers, only to cancel the cards once the rewards were received.
-Given that similar problems exist even in traditional contexts, where users can be easily identified and penalized, it seems that reward farming cannot be solved easily.
-5.2 Decentralized Governance Threats
-Some protocols airdrop governance tokens, as a means to decentralize their governance process.
-However, distributing governance tokens may be dangerous.
-Such tokens allow holders to vote in the protocols’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value [17].
-There is some empirical evidence that suggests that airdrops of governance tokens may be more successful than airdrops of non-governance tokens.
-In a recent analysis of the market capitalization of crypto tokens performed by Makridis et al.
-[53], the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped.
-However, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
-Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system [56, 22, 68].
-Also, to users that do not have the protocol’s best interests in mind, and may even act to the determent of the protocol.
-5.3 Insider Trading
-It is a problematic occurrence within organizations where individuals exploit privileged information to gain a financial advantage over other users of the protocol.
-This practice is widely recognized as a violation of securities laws in traditional financial markets [6], and engaging in such activities can lead to negative responses from the blockchain community [69, 61, 23].
-When someone within a protocol utilizes privileged information to enhance their profits during an airdrop, it can trigger a backlash from the blockchain community.
-Insider traders have advanced knowledge of the metrics used to determine eligibility and the amount of an airdrop for each address.
-This knowledge empowers them to manipulate the system and make a profit.
-For example, there were claims that a head of growth at AltLayer might have used insider information to make a $200000$ USD profit from their airdrop, although later dismissed as a mere coincidence [69].
-However, such incidents can undermine the trust users place in these protocols.
-This issue also raises concerns about fairness, as some users have access to superior and more accurate information than others.
-Identifying these traders is a challenging problem, making it crucial for protocols to provide extensive information to their users.
-In addition, incentives for blockchain data analytics companies and research groups to conduct post-airdrop data audits can help identify insider traders by analyzing address details and transfer patterns [23].
-For this to happen, data availability is key.
-Therefore, protocols should ensure transparency and encourage thorough analysis to maintain integrity and fairness within the blockchain community.
-6 Design Guidelines
-The airdrop design challenges outlined earlier, while indeed sobering, may inform blockchain protocols embarking on the airdrop journey, and can shed light on the path to potential success.
-6.1 Alternative Incentives may Better Sustain User Engagement
-An airdrop’s long-term benefit to a protocol can be tricky to measure, given that the potential benefits may be indirect and hard to quantify, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from protocols to users and is thereafter no longer available for other uses.
-Furthermore, some costs and effects cannot be easily foreseen, like those arising from distributing governance tokens to users.
-Instead of using airdrops, a community could adopt other measures which have a more predictable relation between costs and benefits.
-For example, a very natural alternative to airdrops is for the community to vote to programmatically reward power users with discounts on future interactions.
-When considering L2s, these discounts can be given on transaction fees.
-Thus, if a user receives this reward, it necessarily has to use the protocol again in the future to benefit from the incentive, thus effectively fostering continued user engagement.
-Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the protocol, and the protocol’s costs go solely towards users actively using the system.
-Nevertheless, while fee discounts are more focused towards incentivizing continued interaction within the protocol, they are not without their own challenges.
-For example, one should carefully design a proper discount mechanism, covering both the eligibility criteria for users, and the height of awarded discounts.
-Furthermore, it is not immediately apparent that discounts can attract users to a similar degree as the more immediate and tangible rewards offered by airdrops.
-Another option is to perform multiple airdrops over a relatively long period of time, instead of a single one.
-Although this is still susceptible to some of the pitfalls of standard one-shot airdrops, it may serve to ensure the long-term engagement of the community, and thus avoid the decline in user adoption that some experience immediately after an airdrop.
-A more innovative approach was adopted by Blast, which introduced a point-based reward program [55].
-Under this program, users receive rewards based on the points they accumulate through various activities, such as bridging tokens to the protocol (i.e., transferring users’ funds from another protocol) and participating in a referral program where bringing in more users increases their points.
-Interestingly, Blast received 1.1 billion USD in deposits even before its official launch [43].
-This method provides an evaluative metric for users’ contributions to the protocol, structured around a referral-program model.
-Furthermore, the incorporation of innovative distribution mechanisms demonstrates as a critical role in mitigating the presence of sybils within whitelisted addresses eligible for an airdrop.
-For example, Celestia has proposed a distinctive design that employs GitHub commits as a proxy to assess a user’s contribution to the blockchain ecosystem [27].
-Nevertheless, a potential concern emerges in regard to the possibility of users or farmers engaging in the generation of counterfeit activities on GitHub, aiming to exploit other protocols’ airdrops adopting a strategy analogous to Celestia’s.
-Among other metrics, would be to focus on metrics that are resistant to programmable manipulation, adding a layer of difficulty or cost to automated user accounts.
-6.2 Target Well-known Reputable Entities
-Instead of rewarding pseudonymous users, one can target DAOs that are building relevant apps.
-In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs [24].
-Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students [26] or members of the technical community [25] who wish to take part in research and build tools relevant to the protocol.
-Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world [28].
-6.3 Proactive Oversight and Community Engagement
-While the airdrop is ongoing, the protocol’s data should be carefully analyzed to verify that no one is attempting to game the system.
-A good example was put forth by Linea, whose team uncovered a vulnerability [49] that allowed users to cheat their incentive mechanism.
-The prompt discovery prevented cheaters from obtaining over one-third [^2] A dashboard is available on Dune at https://dune.com/rusty378/guildxyz-pins.
-of NFTsoffered as incentives.
-A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities [71], whether these are actively used or not.
-For example, a disclosure by an AzukiDAO community member allowed the protocol to mitigate a vulnerability relatively quickly [58].
-A protocol’s oversight should extend beyond on-chain data.
-In particular, social media is rife with scammers who advertise “fake” airdrops for their protocols, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds.
-Even protocols that are not planning an airdrop are targeted by such ploys.
-Actively engaging with the community regarding technical matters can prove beneficial, as well.
-For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar [16], who found that it could be implemented more cost-effectively.
-Operating transparently and giving the community visibility into the inner-workings of the protocol can assure that even if technical issues arise, they are met with understanding by users.
-6.4 Rewards Should Scale with Costs
-The ramifications of Goodhart’s law are readily observable in previous airdrops.
-For example, some airdrops attempt to incentivize user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the protocol in the past.
-Consequently, users can game the system by performing meaningless wash-trading transactions [74], meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the protocol.
-This is further compounded by the fact that action-based metrics used by airdrops do not usually account for the actual costs incurred by users for each action.
-For instance, if the number of transactions is used as a metric, one should consider that if transaction fees are low, an airdrop farmer may cheaply perform the required amount of transactions.
-To address this, adopting a quadratic voting system could deter users from artificially inflating transaction volumes, as their rewards would be proportionate to the square root of their activities [11, 48].
-Consequently, the cost to achieve a similar impact would rise, acting as a barrier against manipulation.
-On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop [60].
-To remedy these issues, rewards should scale with the costs that are incurred by users.
-6.5 Data Availability
-Blockchain data is typically publicly accessible but not easily accessible.
-Ensuring data availability is crucial for monitoring airdrop activities and combating Sybil attacks.
-While previous research on airdrops and Sybil detection has primarily focused on single blockchains, our study expands this scope to multiple blockchains.
-Specifically, we collect data from two major Ethereum rollups, Arbitrum and zkSync Era, and gather and label data from significant airdrops, including a major Sybil attack on zkSync Era (the Gemstone “airdrop”).
-Since Sybil detection is increasingly challenging and requires data from multiple chains—given that protocols often reward participants from different chains [27, 4]—data availability is essential for analyzing past airdrops and testing new metrics for future applications.
-However, the rapid growth of blockchain data can make its collection and processing impractical.
-Protocols should ensure their data is available for researchers and companies, helping improve monitoring tools and facilitate learning from the data.
-To support this, we will make our zkSync Era and Ethereum datasets publicly available, enabling others to utilize them for optimizing Sybil detection and airdrop designs.
-7 Related Work
-The majority of recent papers on airdrops concerned themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
-Studies.
-Makridis et.
-al.
-[53] investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges.
-They find that airdrops of governance tokens, in particular, are positively associated with growth in market capitalization and volume.
-Lommers et.
-al.
-[52] offer a broad treatment of the different types of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based).
-They offer examples of projects like Blur and the pitfalls of their airdrops, along with some advice on airdrop implementation strategies.
-In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success, and they provide practical guidance on how to toggle these.
-Fan et.
-al.[21] study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristics of community members and the effectiveness of airdrop.
-They find that users who receive more rewards tend to take positive actions towards the community.
-They summarize several arbitrage patterns and confirm that the current detection is not sufficient to detect airdrop hunters.
-They use an unsupervised hierarchical clustering method to capture the behaviors of eligible airdrop addresses after receiving rewards and to divide them into interpretable categories.
-Other works focused on addressing the problem of Sybil detection, providing valuable insights into identifying airdrop farmers.
-Commonly adopted practices include graph network analysis [21, 51] and machine learning techniques [36, 79, 73, 13, 59].
-The former involves identifying prevalent transfer patterns, such as sequences, radial structures, or a combination of both.
-The latter utilizes approaches like node embedding and employs universal pre-trained transformer encoders like BERT for effective fraud detection [36] or clustering [73, 59].
-Allen [2] examines airdrop design as co-evolutionary process and makes three contributions.
-First, applying a co-evolutionary lens to airdrops, where participants learn and adapt their strategies (e.g., new criteria, on-chain farming) both within and between web3 ecosystems.
-Second, nine airdrop cases studies (e.g., Optimism, Arbitrum, Blur) that demonstrate the expanding set of airdrop strategies (e.g., task-based claiming).
-Third, implications for airdrop design including more feedback loops to facilitate learning (e.g., multi-round airdrops), ongoing innovation in airdrop criteria (e.g., randomness), and a reversion back to simple designs for some projects (because of the costs of airdrop design).
-Allen et.
-al.[3] consider why projects might choose this unusual and costly means of token distribution.
-It considers a selection of high-profile airdrops as case studies between 2014 and 2022.
-This is the first comprehensive analysis of the rationales and mechanisms of Web3 token airdrops.
-They find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralization of ownership and control of a project (building community, providing regulatory protection, and enhancing security).
-While many commentators describe airdrops as primarily for marketing, they claim that marketing is a weak rationale for airdropping tokens and the evidence of successful marketing-driven airdrops lacking.
-Technical Aspects of Airdrops  Frowis et.
-al.
-[31] point out that efficient transfers to a large number of recipients in the case of airdrops present problems on Ethereum.
-They find that cost savings of a factor of two are in fact possible, but require specific provisions in the smart contract implementing the token system.
-Pull-based approaches, which use on-chain interaction with recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient.
-Total costs are broadly linear in the number of recipients, regardless of the technical approach.
-Wahby et.
-al.
-[75] point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities.
-In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely used user credentials (ECDSA and RSA).
-Private airdrop for RSA builds upon a new zero-knowledge argument of knowledge of the factorization of a committed secret integer, which may be of independent interest.
-Their fastest implementation takes 40–180 ms to generate and 3.7–10 ms to verify an RSA private airdrop signature.
-8 Conclusions
-In this study, we investigated the prevalent issues with airdrops and developed guidelines to improve their impact.
-Our study analyzed six notable airdrops, uncovering a pattern in which recipients quickly traded their airdropped tokens on exchanges.
-Specifically, 36.51% of ENS tokens, 35.35% of dYdX tokens, and 53.47% of 1inch tokens were exchanged soon after distribution.
-Generally, these tokens were sold after only 1 or 2 transactions, with a median of 2 transactions, demonstrating that airdrops do not promote sustained engagement or attract contributors of significant value.
-Furthermore, our examination of Arbitrum indicated a significant increase in total daily fees during the airdrop event, but a decrease in transaction count per address afterward.
-In particular, other protocols outperformed Arbitrum in the absence of an airdrop.
-The average transaction fees across all protocols followed a similar trend since June 2023, suggesting that airdrops may not be the primary driver of user base growth and can impact activity on other chains.
-We also acknowledge the common airdrop challenges related to airdrop farmers, governance token distribution, and insider trading.
-These challenges serve as valuable lessons for blockchain protocols considering airdrop tokens and offer insight into the potential path to success.
-As a measure to avoid these pitfalls, we propose design guidelines for airdrop campaigns, which are summarized below.
-Alternative incentives may better sustain user engagement.
-An interesting approach involves offering alternative incentives, such as fee discounts for future interactions within the protocol, to foster long-term engagement.
-Additionally, implementing multi-round airdrops allows for learning from past rounds, enabling iterative improvements to address previously identified issues.
-Another approach involves exploring innovative metrics to assess the eligibility of an address for receiving an airdrop.
-This might also be a promising tactic to address concerns related to airdrop farming who typically tries to game the system by utilizing past airdrop actions.
-However, this approach may not completely solve the challenge of insider trading.
-Target well-known reputable entities.
-Prioritizing established and reputable entities, including those building on the protocol, research groups, tech communities, and students, can contribute to sustained engagement.
-Funding these entities or directing airdrops towards DAOs developing relevant dApps presents an opportunity to attract value-driven users and promote long-term involvement.
-Proactive oversight and community engagement.
-Critical to the success of any airdrop is proactive oversight and community engagement.
-Analyzing data from past airdrops, even those on different chains, provides insights into pitfalls and allows for adjustments in future airdrop strategies.
-Open communication channels with the community help understand their interests and desired protocol functionalities.
-Rewards should scale with costs.
-Ensure that rewards scale proportionally with associated costs to prevent exploitation by airdrop farmers.
-Aligning rewards with costs is a strategic measure to mitigate potential gaming of the system.
-In conclusion, our paper offers insights into airdrop challenges and proposes practical guidelines for designing effective airdrop campaigns.
-We hope these suggestions inspire blockchain protocols, researchers, and practitioners.
-Finally, in our commitment to promoting scientific reproducibility, we plan to share the dataset and scripts utilized in this study [57].